### v2

# latexfontFont shape

Abstract
Based on the recently proposed framework of general relativistic stochastic mechanics [1, 2] and stochastic thermodynamics [3] at the ensemble level, this work focuses on general relativistic stochastic thermodynamics at the trajectory level. The first law of stochastic thermodynamics is reformulated and the fluctuation theorems are proved on this level, with emphasis on maintaining fully general covariance and on the choice of observers.
Keywords: Langevin equation, fluctuation theorem, time-reversal symmetry, general relativity
1 Introduction
One of the central problem in modern statistical physics is the origin of irreversibility in macroscopic and mesoscopic systems. This problem can be traced back to Boltzmann’s efforts in proving the second law of thermodynamics starting from deterministic mechanics, which has resulted in the well-known H-theorem [4]. However, the debates about the validity of (the assumptions of) the H-theorem have lasted for more than a century. The most acute critics to the H-theorem is reflected by Loschmidt paradox [5, 6], which roughly states that the macroscopic arrow of time cannot possibly arise from the underlying microscopic mechanics obeying time reversal symmetry (TRS). This paradox remains unresolved until the 1990s, when numerous works [7, 8, 9] emerged, revealing that the forward and reversed processes are not probabilistically equally likely, provided certain dissipative effect exists on the mechanical level. These results, known as fluctuation theorems, largely resolved the debates and paradoxes related to the H-theorem. In particular, the most questioned molecular chaos hypothesis adopted in proving the H-theorem is completely avoided in proving the fluctuation theorems.
Almost at the same time, Sekimoto [10] established the first law of stochastic thermodynamics on the trajectory level by use of the Langevin equation. This formulation establishes a connection between stochastic mechanics and fluctuation theorems. In 2005, Seifert [11] presented a version of fluctuation theorem based on the overdamped Langevin equation. Subsequently, several fluctuation theorems based on non-relativistic stochastic mechanics were presented [12, 13, 14, 15], making it clear that stochastic mechanics provides an ideal starting point for constructing fluctuation theorems and interpreting the origin of irreversibility.
Most of the works mentioned above are carried out in the non-relativistic regime. Nowadays, it is widely acknowledged that thermodynamics and relativity are both concentrated on the universal principles which every physical system must obey. It is important to establish fluctuation theorems based on these universal principle theories. However, since the spacetime symmetry in relativity imposes stronger protection to the time reversal invariance, the extension of fluctuation theorems to the relativistic regime proves to be more difficult. The central difficulty lies in how to incorporate the breaking of time reversal invariance while still maintaining relativistic covariance. Refs. [16, 17, 18, 19, 20, 21] considered the extension of fluctuation theorems to the special relativistic regime. However, the long awaited general relativistic extension is still beyond our human’s sight.
Recently, we developed a framework for dealing with stochastic mechanics on curved spacetime and investigated the general relativistic stochastic thermodynamics based on this framework [1, 2, 3]. Meanwhile, we also established a version of fluctuation theorem on an arbitrary curved Riemannian manifold [15]. The aim of the present work is to employ the framework established in [1, 2, 3] and make use of the technique introduced in [15] to construct a version of fluctuation theorem based on a fully general relativistic description of stochastic mechanics and stochastic thermodynamics.
In our framework of relativistic stochastic mechanics [1, 2], it is important to liberate the observer from the coordinate system. Rather than fixing the zeroth component of the coordinate system, we utilize the observer’s proper time $t$ to to label the configuration space $\mathcal{S}_{t}$ and the space of microstates $\Sigma_{t}$, enabling our theory to possess general covariance. For this purpose, we first clarify the geometry of the space of $\mathcal{S}_{t}$ and $\Sigma_{t}$ in Sec. 2. To make the construction more self-contained, we provide a brief review of the basics of our framework of relativistic stochastic mechanics and relativistic stochastic thermodynamics in Sec. 3, and the first law of relativistic stochastic thermodynamics at trajectory level is established in this section also. The separation of the observer from the coordinate system is also important in describing the time reversal transformation (TRT) in curved spacetime. In contrast to the usual practice in special relativistic theories (including special relativistic field theories) in which the TRT is often described as a coordinate transformation $(t,x^{i})\mapsto(-t,x^{i})$, our general covariant framework calls for interpreting the TRT as a transformation from a future-directed observer to a past-directed observer. This concept is elucidated in detail in Sec. 4. Ref.[1] presents a method to establish the covariant relativistic Langevin equation from the perspective of the observer. Based on this approach, the forward and reversed processes are respectively defined as stochastic processes from the perspectives of the future-directed and past-directed observers. In Sec. 5, we demonstrate that the forward and reversed processes possess distinct probabilities, thus breaking the TRS and leading to a version of fluctuation theorem on curved spacetime manifold. Finally, in Sec. 6, we provide brief concluding remarks.
We maintain the notations and conventions consistent with the previous works [1, 2]. To distinguish random variables from their realizations, the former are labeled with extra tildes. For instance, $\tilde{x}$ represents a random variable, while $x$ denotes its realization. Several manifolds of different dimensions will be relevant in our discussion. These include the $(d+1)$-dimensional spacetime manifold $\mathcal{M}$ with metric $g_{\mu\nu}(x)$ of signature (,+,,+)(-,+$\cdots$,+)( - , + , ⋯ , + ), its tangent bundle $T\mathcal{M}$ of dimension $(2d+2)$, and certain submanifolds within the tangent bundle. To distinguish tensors on these different manifolds, we introduce distinct indices. Lower-case Greek letters such as $\alpha,\beta,\mu,\nu,\rho,\ldots$, are used as concrete indices and range from 00 to $d$. Lower-case Latin letters like $i,j,k,l,m,\ldots$ are also used as concrete indices, which range from $1$ to $d$. Lastly, lower-case Latin letters $a,b,c,\ldots$ are used as abstract indices. This paper is intended to be as mathematically rigorous as possible. A more concise summary of the main results can be found in Ref.[22].
2 Geometry of the space of microstates
Statistical physics is built on top of the space of microstates. For systems consisting of classical massive particles, the space of microstates can be subdivided into the configuration space and the momentum space. In relativistic context, the configuration space is a subspace of the spacetime manifold $\mathcal{M}$ consisting of simultaneous events at a given instance of time, while the momentum space for each individual particle is a subspace of the tangent (or cotangent) space of the spacetime at a given event which obeys the mass shell condition
$\mathcal{H}(x,p):=g_{\mu\nu}(x)p^{\mu}p^{\nu}+m^{2}=0.$ (1)
Please be reminded that, due to the non-degeneracy of the spacetime metric $g_{\mu\nu}(x)$, the tangent and cotangent spaces are dual to each other and both are equally well for describing the momentum space. In this work, we adopt the tangent space description. When cotangent space variables appear, they are considered to be linear functions of the tangent space variables, e.g. $p_{\mu}=g_{\mu\nu}(x)p^{\nu}$.
In principle, the space of microstates for a single particle should be considered as a submanifold of the mass shell bundle
$\Gamma_{m}:=\{(x,p)\in T\mathcal{M}|\mathcal{H}(x,p)=0\},$ (2)
with the configuration space taken to be a subspace of the spacetime manifold $\mathcal{M}$ consisting of simultaneous events. To clarify the concept of simultaneous events, we need to introduce an arbitrary observer field which is encoded by a normalized timelike vector field $Z^{\mu}$ obeying $g_{\mu\nu}Z^{\mu}Z^{\nu}=-1$. For convenience, we shall refer to this arbitrary observer field as Alice. When considering the motion of a Brownian particle inside a heat reservoir, there is a particular observer field that comoves with the reservoir. This particular observer field will be referred to as Bob.
Consider the worldline $x_{\tau}$ of a massive relativistic particle of which $\tau$ is its proper time. If the time orientations of Alice and the particle align, i.e. $g_{\mu\nu}p^{\mu}_{\tau}Z^{\nu}<0$, where $p^{\mu}_{\tau}:=m{\rm d}x^{\mu}_{\tau}/{\rm d}\tau$, the part of the mass shell bundle in which the phase trajectory lies is defined as the future mass shell bundle relative to Alice and is denoted as $\Gamma^{+}_{m}$,
$\Gamma^{+}_{m}:=\{(x,p)\in T\mathcal{M}|\mathcal{H}(x,p)=0,\ g_{\mu\nu}p^{\mu}Z^{\nu}<0\}.$ (3)
For notational convenience, the phase trajectory is denoted as $X_{\tau}=(x_{\tau},p_{\tau})$, which is the uplift of the particle’s worldline $x_{\tau}$ into the bundle $\Gamma^{+}_{m}$.
Figure 1: The worldline of a particle is lifted to different regions of the mass shell bundle by different observers.
If we consider another observer field with proper velocity $C^{\mu}=-Z^{\mu}$, referred to as Carol, whose time orientation is opposite to that of Alice, the definition of the future mass shell bundle relative to Carol will be opposite to that relative to Alice. To avoid confusion, we will designate Alice as the future-directed observer and Carol as the past-directed observer, and throughout this paper, the future and past mass shell bundles are always defined relative to Alice. This designation is arbitrary because, in the presence of TRS, the future- and the past-directed observer fields are indistinguishable on the level of deterministic mechanics.
The future mass shell bundle is also a fiber bundle based on $\mathcal{M}$. Its fiber, denoted by $(\Gamma^{+}_{m})_{x}$, is the momentum space of the relativistic particle. We use the calligraphy fonts, like $\mathcal{F}$, $\mathcal{R}$ and $\mathcal{K}$, to denote tensors on $(\Gamma_{m}^{+})_{x}$, and the cursive fonts, like $\mathscr{N}$, $\mathscr{Z}$ and $\mathscr{L}$, to denote tensors on $\Gamma_{m}^{+}$.
The future mass shell bundle still is not the space of microstates, because the base manifold $\mathcal{M}$ is not the configuration space. There are different ways to drop the temporal dimension of $\mathcal{M}$, e.g. either by fixing the zeroth component $x^{0}$ of coordinates or by fixing the proper time $t$ of Alice. The first approach lacks apparent general covariance. The core idea of our framework for relativistic stochastic mechanics is to keep the manifest general covariance and emphasize the role of observer choice. The details can be found in Ref.[1]. Here we list some of the key points:
  1. 1.
    The proper time $t$ of Alice can be extended into a scalar field $t(x)$ on $\mathcal{M}$;
  2. 2.
    The configuration space relative to Alice is defined as a constant proper time slice of $\mathcal{M}$, i.e. a spacelike hypersurface $\mathcal{S}_{t}:=\{x\in\mathcal{M}|t(x)=t\}$ in $\mathcal{M}$. Moreover, the proper velocity field $Z^{\mu}$ of Alice is the unit normal vector field of $\mathcal{S}_{t}$;
  3. 3.
    The space of microstates is defined as a hypersurface in $\Gamma^{+}_{m}$ as a constant time slice, i.e. $\Sigma_{t}^{+}:=\{(x,p)\in\Gamma^{+}_{m}|t(x)=t\}$;
  4. 4.
    Since the gradient of $t$ must be a normal co-vector of $\mathcal{S}_{t}$, $Z_{\mu}$ must be collinear with $\nabla_{\mu}t$. Denoting $|\nabla_{\mu}t|$ as $\lambda$, we have $\partial_{\mu}t=-\lambda Z_{\mu}$.
Since $Z$ is the unit future-directed normal vector field of $\mathcal{S}_{t}$, the volume element on $\mathcal{S}_{t}$ can be written as the interior derivative of the volume element on $\mathcal{M}$ along $Z^{\mu}$:
$\eta_{\mathcal{S}_{t}}:=\iota_{Z}\eta_{\mathcal{M}},$ (4)
where $\iota$ represents interior derivative, and $\eta_{\mathcal{M}}=g^{1/2}{\rm d}x^{0}\wedge{\rm d}x^{1}\wedge\cdots\wedge{\rm d}x^{d}$ is the volume element on $\mathcal{M}$, with $g=|{\rm det}(g_{\mu\nu})|$. Since $\iota_{Z}({\rm d}x^{\mu})=Z^{\mu}$ and ${\rm d}x^{0}=-\partial_{i}t{\rm d}x^{i}/\partial_{0}t$ on $\mathcal{S}_{t}$, we have
$\eta_{\mathcal{S}_{t}}$ $=g^{1/2}\iota_{Z}({\rm d}x^{0}\wedge{\rm d}x^{1}\wedge\cdots\wedge{\rm d}x^{d})$
$=g^{1/2}\sum_{\rho=0}^{d}(-1)^{\rho}\iota_{Z}({\rm d}x^{\rho}){\rm d}x^{0}\wedge\cdots\wedge{\rm d}x^{\rho-1}\wedge{\rm d}x^{\rho+1}\wedge\cdots\wedge{\rm d}x^{d}$
$=g^{1/2}Z^{0}{\rm d}x^{1}\wedge\cdots\wedge{\rm d}x^{d}-g^{1/2}\sum_{i=1}^{d}Z^{i}{\rm d}x^{1}\wedge\cdots\wedge{\rm d}x^{0}\wedge\cdots\wedge{\rm d}x^{d}$
$=g^{1/2}\left[Z^{0}+\frac{1}{\partial_{0}t}\partial_{i}tZ^{i}\right]{\rm d}x^{1}\wedge\cdots\wedge{\rm d}x^{d}$
$=-\frac{\lambda g^{1/2}}{\partial_{0}t}{\rm d}x^{1}\wedge\cdots\wedge{\rm d}x^{d}.$ (5)
We shall also need to make use of the volume elements on $\Gamma^{+}_{m}$ and $\Sigma_{t}^{+}$. Since both of them are submanifolds of $T\mathcal{M}$, it is appropriate to begin from the geometry of $T\mathcal{M}$. The non-degenerate metric on $T\mathcal{M}$ is known as the Sasaki metric [23] $\hat{g}_{ab}$, which is determined by the metric $g_{\mu\nu}$ of the base manifold $\mathcal{M}$,
$\hat{g}_{ab}:=g_{\mu\nu}{\rm d}x^{\mu}{}_{a}{\rm d}x^{\nu}{}_{b}+g_{\mu\nu}\theta^{\mu}{}_{a}\theta^{\nu}{}_{b},\quad\hat{g}^{ab}:=g^{\mu\nu}e_{\mu}{}^{a}e_{\nu}{}^{b}+g^{\mu\nu}\left(\frac{\partial}{\partial p^{\mu}}\right)^{a}\left(\frac{\partial}{\partial p^{\nu}}\right)^{b},$ (6)
where
$\theta^{\mu}:={\rm d}p^{\mu}+\varGamma^{\mu}{}_{\alpha\beta}p^{\alpha}{\rm d}x^{\beta},\quad e_{\mu}:=\frac{\partial}{\partial x^{\mu}}-\varGamma^{\alpha}{}_{\mu\beta}p^{\beta}\frac{\partial}{\partial p^{\alpha}},$
and $\varGamma^{\mu}{}_{\alpha\beta}$ is the Christoffel connection associated with $g_{\mu\nu}$. The corresponding volume element reads
$\eta_{T\mathcal{M}}$ $=g~{}{\rm d}x^{0}\wedge{\rm d}x^{1}\wedge\cdots\wedge{\rm d}x^{d}\wedge\theta^{0}\wedge\cdots\wedge\theta^{d}$
$=g~{}{\rm d}x^{0}\wedge{\rm d}x^{1}\wedge\cdots\wedge{\rm d}x^{d}\wedge{\rm d}p^{0}\wedge\cdots\wedge{\rm d}p^{d}.$ (7)
As a hypersurface in $T\mathcal{M}$, $\Gamma^{+}_{m}$ has the unit normal (co)vector
$\hat{N}_{a}=\frac{1}{|{\rm d}\mathcal{H}|}{\rm d}\mathcal{H}_{a}=\frac{p_{\mu}}{m}\theta^{\mu}{}_{a},\qquad\hat{N}^{a}=\hat{g}^{ab}\hat{N}_{b}=\frac{p^{\mu}}{m}\left(\frac{\partial}{\partial p^{\mu}}\right)^{a},$ (8)
giving rise to the induced metric
$\hat{h}_{ab}$ $:=\hat{g}_{ab}+\hat{N}_{a}\hat{N}_{b}=g_{\mu\nu}{\rm d}x^{\mu}{}_{a}{\rm d}x^{\nu}{}_{b}+\Delta_{\mu\nu}(p)\theta^{\mu}{}_{a}\theta^{\nu}{}_{b},$ (9)
$\Delta_{\mu\nu}(p)$ $=g_{\mu\nu}+\frac{1}{m^{2}}p_{\mu}p_{\nu},$ (10)
where the second term is the induced metric on $(\Gamma^{+}_{m})_{x}$:
$h_{ab}:=\Delta_{\mu\nu}(p)\theta^{\mu}{}_{a}\theta^{\nu}{}_{b}.$ (11)
Consequently, we get the volume element
$\eta_{\Gamma^{+}_{m}}$ $:=\iota_{\hat{N}}\eta_{T\mathcal{M}}=-\frac{m}{p_{0}}g~{}{\rm d}x^{0}\wedge\cdots\wedge{\rm d}x^{d}\wedge{\rm d}p^{1}\wedge\cdots\wedge{\rm d}p^{d}.$ (12)
It is easy to see that the volume element $\eta_{\Gamma^{+}_{m}}$ can be factorized into the wedge product of $\eta_{\mathcal{M}}$ and $\eta_{(\Gamma^{+}_{m})_{x}}$,
$\eta_{\Gamma^{+}_{m}}=\eta_{\mathcal{M}}\wedge\eta_{(\Gamma_{m}^{+})_{x}},$ (13)
where
$\eta_{(\Gamma^{+}_{m})_{x}}:=-\frac{m}{p_{0}}g^{1/2}{\rm d}p^{1}\wedge\cdots\wedge{\rm d}p^{d}$ (14)
is the volume element on the fiber space $(\Gamma^{+}_{m})_{x}$. Finally, since $\Sigma^{+}_{t}=\bigcup_{x\in\mathcal{S}_{t}}(\Gamma_{m}^{+})_{x}$, the volume element on $\Sigma^{+}_{t}$ also has a factorized form,
$\eta_{\Sigma^{+}_{t}}=\eta_{\mathcal{S}_{t}}\wedge\eta_{(\Gamma_{m}^{+})_{x}}=\frac{m\lambda g}{p_{0}\partial_{0}t}{\rm d}x^{1}\wedge\cdots{\rm d}x^{d}\wedge{\rm d}p^{1}\wedge\cdots\wedge{\rm d}p^{d}.$ (15)
3 General relativistic Langevin systems
This section is intended for a brief review of the framework for general relativistic stochastic mechanics and thermodynamics [1, 2, 3] in order to fix the notations and make the forthcoming presentation for the proof of fluctuation theorem self-contained.
3.1 Covariant Langevin equations
Let us consider a relativistic Brownian particle carrying an electric charge $q$ and moving in a heat reservoir residing in the curved spacetime $\mathcal{M}$ and subjecting to an external electromagnetic field $F=F_{\mu\nu}{\rm d}x^{\mu}\wedge{\rm d}x^{\nu}$. A version of the corresponding general covariant Langevin equation (referred to as LEτ) that takes the proper time $\tau$ of the particle as evolution parameter reads
${\rm d}\tilde{x}_{\tau}^{\mu}$ $=\frac{\tilde{p}_{\tau}^{\mu}}{m}{\rm d}\tau,$ (16)
${\rm d}\tilde{p}_{\tau}^{\mu}$ $=\xi^{\mu}_{\tau}{\rm d}\tau+\mathcal{F}^{\mu}_{\text{dp}}{\rm d}\tau+\mathcal{F}^{\mu}_{\text{em}}{\rm d}\tau-\frac{1}{m}\varGamma^{\mu}{}_{\alpha\beta}\tilde{p}_{\tau}^{\alpha}\tilde{p}_{\tau}^{\beta}{\rm d}\tau,$ (17)
where $\mathcal{F}^{\mu}_{\text{dp}}:=\mathcal{K}^{\mu\nu}U_{\nu}$ is the damping force with $\mathcal{K}^{\mu\nu}$ being the damping coefficient which transforms as a tensor under general coordinate transformations, $U^{\mu}$ is the proper velocity of the heat reservoir, $\mathcal{F}^{\mu}_{\text{em}}:=\frac{q}{m}F^{\mu}{}_{\nu}\tilde{p}^{\nu}_{\tau}$ is the electromagnetic force, and
$\xi_{\tau}^{\mu}:=\mathcal{R}^{\mu}{}_{\mathfrak{a}}\circ_{S}{\rm d}\tilde{w}^{\mathfrak{a}}_{\tau}/{\rm d}\tau+\mathcal{F}_{\text{add}}^{\mu}$ (18)
is the stochastic force, which is consisted of a random force term $\mathcal{R}^{\mu}{}_{\mathfrak{a}}\circ_{S}{\rm d}\tilde{w}^{\mathfrak{a}}_{\tau}/{\rm d}\tau$ encoding the Stratonovich type coupling between the stochastic amplitudes $\mathcal{R}^{\mu}{}_{\mathfrak{a}}$ with a set of independent Gaussian noises ${\rm d}\tilde{w}^{\mathfrak{a}}_{\tau}$ obeying the probability distribution
$\Pr[{\rm d}\tilde{w}^{\mathfrak{a}}_{\tau}={\rm d}w^{\mathfrak{a}}]=\frac{1}{(2\pi{\rm d}\tau)^{d/2}}\exp\left(-\frac{1}{2}\frac{\delta_{\mathfrak{a}\mathfrak{b}}{\rm d}w^{\mathfrak{a}}{\rm d}w^{\mathfrak{b}}}{{\rm d}\tau}\right),$ (19)
and an additional stochastic force term
$\mathcal{F}_{\text{add}}^{\mu}=\frac{\delta^{\mathfrak{a}\mathfrak{b}}}{2}\mathcal{R}^{\mu}{}_{\mathfrak{a}}\nabla^{(h)}_{i}\mathcal{R}^{i}{}_{\mathfrak{b}}$ (20)
in which $\nabla^{(h)}_{i}$ denotes the spatial components of the covariant derivative associated to the metric (11) on the momentum space. Both the random force and the additional stochastic force terms arise from the interaction of the Brownian particle with the heat reservoir. The additional stochastic force is required in order for the Brownian particle to be able to reach equilibrium in the long time limit [24]. Each component of $\mathcal{R}^{\mu}{}_{\mathfrak{a}}$ is assumed to be smoothly dependent on the coordinates on $\Sigma_{t}^{+}$, and for each fixed $\mathfrak{a}=1,2,\cdots d$, $\mathcal{R}^{\mu}{}_{\mathfrak{a}}$ transforms as a vector under general coordinate transformations. As did in Refs. [1, 2], we use tilded and un-tilded symbols to denote the random variables and their realizations.
Although LEτ is perfectly general covariant and encodes all necessary factors that affect the motion of the Brownian particle, there is still some drawbacks which call for an alternative version of the covariant Langevin equation. The problem is connected to the choice of evolution parameter $\tau$. Since
${\rm d}t=\partial_{\mu}t{\rm d}\tilde{x}^{\mu}=-\lambda Z_{\mu}{\rm d}\tilde{x}^{\mu}=-\lambda Z_{\mu}\frac{{\rm d}\tilde{x}^{\mu}}{{\rm d}\tau}{\rm d}\tau=-\lambda\frac{Z_{\mu}\tilde{p}^{\mu}}{m}{\rm d}\tau=\gamma(\tilde{x},\tilde{p}){\rm d}\tau,$ (21)
we have ${\rm d}\tau=\gamma^{-1}(\tilde{x},\tilde{p}){\rm d}t$. Therefore, from the perspective of the observer Alice, the proper time $\tau$ of the Brownian particle becomes a random variable. To avoid this inconvenience, a reparametrization scheme is adopted in [1], with
$\tilde{X}_{\tau}=(\tilde{x}_{\tau},\tilde{p}_{\tau})~{}\mapsto~{}\tilde{Y}_{t}=(\tilde{y}_{t},\tilde{k}_{t}),$
where
$\tilde{y}_{t}:=\tilde{x}_{\tilde{\tau}_{t}},\quad\tilde{k}_{t}:=\tilde{p}_{\tilde{\tau}_{t}}.$
This leads to the following alternative version of covariant Langevin equation which is referred to as LEt for short,
${\rm d}\tilde{y}_{t}^{\mu}$ $=\frac{\tilde{k}^{\mu}_{t}}{m}\gamma^{-1}{\rm d}t,$ (22)
${\rm d}\tilde{k}_{t}^{\mu}$ $=\hat{\xi}_{t}^{\mu}\gamma^{-1}{\rm d}t+\mathcal{F}^{\mu}_{\text{dp}}\gamma^{-1}{\rm d}t+\mathcal{F}^{\mu}_{\text{em}}\gamma^{-1}{\rm d}t-\frac{1}{m}\varGamma^{\mu}{}_{\alpha\beta}\tilde{k}^{\alpha}_{t}\tilde{k}^{\beta}_{t}\gamma^{-1}{\rm d}t.$ (23)
The new stochastic force $\hat{\xi}^{\mu}_{t}$ reads
$\hat{\xi}^{\mu}_{t}:=\gamma^{1/2}\mathcal{R}^{\mu}{}_{\mathfrak{a}}\circ_{S}{\rm d}\tilde{W}_{t}^{\mathfrak{a}}/{\rm d}t+\mathcal{F}^{\mu}_{\text{add}}-\frac{1}{2}\mathcal{D}^{\mu i}\gamma^{1/2}\nabla^{(h)}_{i}\gamma^{-1/2},$ (24)
in which
${\rm d}\tilde{W}^{\mathfrak{a}}_{t}=\gamma^{1/2}(\tilde{Y}_{t}){\rm d}\tilde{w}^{\mathfrak{a}}_{\tilde{\tau}_{t}}$
are still Gaussian noises but with the variance changed from ${\rm d}\tau$ to ${\rm d}t$, and $\mathcal{D}^{\mu\nu}:=\mathcal{R}^{\mu}{}_{\mathfrak{a}}\delta^{\mathfrak{a}\mathfrak{b}}\mathcal{R}^{\nu}{}_{\mathfrak{b}}$ is the diffusion tensor.
In this work, we assume that the diffusion tensor has rank $d$, so that $\mathcal{D}^{ij}$ is a full-rank symmetric matrix. This also requires that $\mathcal{R}^{i}{}_{\mathfrak{a}}$ is a $d\times d$ full-rank matrix. The above assumption is necessary and sufficient to ensure that the Brownian particle couples to the heat reservoir in every spatial direction. Please keep in mind that the choice for the stochastic amplitudes $\mathcal{R}^{\mu}{}_{\mathfrak{a}}$ is non-unique. Different choices correspond to different Langevin systems. The result of the present work does not require the explicit values for the stochastic amplitudes and should be valid for any choices obeying the above assumption.
3.2 Reduced Fokker-Planck equation
Using the diffusion operator method [27], the reduced Fokker-Planck equation (RFPE) associated with LEτ or LEt is obtained in Ref. [2],
$\dfrac{1}{m}\mathscr{L}_{F}(\varphi)=\nabla_{i}^{(h)}\mathcal{I}^{i}[\varphi],$ (25)
where
$\mathscr{L}_{F}:=p^{\mu}e_{\mu}+qF^{\mu}{}_{\nu}p^{\nu}\frac{\partial}{\partial p^{\mu}}$ (26)
is the Liouville vector field for a charged particle,
$\mathcal{I}[\varphi]=\left[\frac{1}{2}\mathcal{D}^{ij}\nabla^{(h)}_{j}\varphi-\mathcal{K}^{i\nu}U_{\nu}\varphi\right]\frac{\partial}{\partial\breve{p}^{i}}$ (27)
is a vector field which is connected to the heat transfer rate from the heat reservoir to the Brownian particle via [1, 3]
$Q[\varphi]:=\int_{(\Gamma^{+}_{m})_{x}}\eta_{(\Gamma^{+}_{m})_{x}}Z_{\nu}\mathcal{I}^{\nu}[\varphi],$ (28)
and the definition for the derivative operator $\frac{\partial}{\partial\breve{p}^{i}}$ is provided in Appendix A. The round and square brackets around $\varphi$ have different meanings: $\mathscr{L}_{F}(\varphi)$ represents the action of the vector field $\mathscr{L}_{F}$ on the scalar $\varphi$, while $\mathcal{I}^{i}[\varphi]$ implies that the vector field $\mathcal{I}^{i}$ is dependent on $\varphi$. Such convention will be used throughout this paper.
It is important to point out that the one particle distribution function (1PDF) $\varphi$ appearing in the RFPE is not a probability distribution in $\Sigma^{+}_{m}$. To see this, we recall that the probability current associated with the RFPE (25) is
$\mathscr{J}[\varphi]=\frac{\varphi}{m}\mathscr{L}_{F}-\mathcal{I}[\varphi].$ (29)
Therefore, the probability distribution function on $\Sigma^{+}_{t}$ should be
$f:=-{\mathscr{Z}_{a}}\mathscr{J}^{a}[\varphi]=-\frac{1}{m}p^{\mu}Z_{\mu}\varphi=\gamma\lambda^{-1}\varphi,$ (30)
where $\mathscr{Z}=Z^{\mu}e_{\mu}$ is the unit normal vector of $\Sigma_{t}^{+}$.
Since $\mathcal{I}[\varphi]$ is proportional to the heat transfer rate, the condition for the Brownian particle to reach detailed thermal equilibrium with the reservoir is $\mathcal{I}[\varphi_{\text{eq}}]=0$, which yields the equilibrium distribution
$\varphi_{\text{eq}}=\operatorname{e}^{-\alpha+\beta_{\mu}p^{\mu}},\quad\beta_{\mu}:=\beta U_{\mu},$ (31)
provided that the covariant Einstein relation
$\mathcal{D}^{\mu\nu}=2\beta^{-1}\mathcal{K}^{\mu\nu}$ (32)
holds and that $\alpha$ and $\beta_{\mu}$ obey the following equations
$\nabla_{\mu}\alpha+q\beta^{\nu}F_{\mu\nu}=0,\qquad\nabla_{(\mu}\beta_{\nu)}=0,$ (33)
which are simple consequences of the Liouville equation $\mathscr{L}_{F}(\varphi_{\text{eq}})=0$. Eq. (33) implies that $\beta_{\mu}$ is a Killing vector field, while the relation $\beta_{\mu}=\beta U_{\mu}$ implies that it is timelike. Therefore, $\beta_{\mu}$ must be timelike Killing. This leads to the conclusion that the existence of the equilibrium distribution (31) requires the spacetime to be at least stationary.
As we have argued in Ref. [2], the equilibrium state is intrinsic to the system which is independent of the choice of observer. However, the parameters that characterize the equilibrium state is indeed observer-dependent. It has been shown [29, 2] that the parameters $\alpha$ and $\beta$ are related to the chemical potential and the temperature observed by Bob via
$\beta=\frac{1}{T_{\rm B}},\qquad\alpha=-\frac{\mu_{\rm B}}{T_{\rm B}}.$ (34)
The equilibrium distribution (31) is recognized to be precisely the Jütnner distribution which is also obeyed by particles of the heat reservoir.
3.3 Thermodynamic relations
The definition of the energy of a charged relativistic particle is non-unique. For instance, both the kinematic momentum $p^{\mu}$ and canonical momentum $P^{\mu}:=p^{\mu}+qA^{\mu}$ can be used for defining the energy [25]:
$E:=-Z_{\mu}p^{\mu},\qquad H:=-Z_{\mu}P^{\mu}.$ (35)
As an analogy of non-relativistic case, $E$ can be viewed as kinematic energy and $H$ can be viewed as the sum of kinematic energy and electromagnetic potential energy $-qZ_{\mu}A^{\mu}$. The non-uniqueness for the definition of energy also appeared in the non-relativistic stochastic thermodynamics [15]. However, such non-uniqueness does not affect the description of heat in the first law of stochastic thermodynamics. We will show that the same situation also occurs the relativistic case.
Since the microstate of the Brownian particle is described by a set of random variables, the energy of the Brownian particle also depends on the same set of random variables,
$\tilde{E}_{\tau}:=E(\tilde{x}_{\tau},\tilde{p}_{\tau}),\qquad\tilde{H}_{\tau}:=H(\tilde{x}_{\tau},\tilde{p}_{\tau}).$ (36)
Since LEτ is a system of Stratonovich-type stochastic differential equations, the chain rule is available
${\rm d}\tilde{E}_{\tau}$ $=\frac{\partial E}{\partial x^{\mu}}{\rm d}\tilde{x}^{\mu}_{\tau}+\frac{\partial E}{\partial p^{\mu}}{\rm d}\tilde{p}^{\mu}_{\tau}$
$=-Z_{\mu}\left[\xi_{\tau}^{\mu}+\mathcal{F}_{\text{dp}}^{\mu}\right]{\rm d}\tau-\frac{\tilde{p}_{\tau}^{\mu}\tilde{p}_{\tau}^{\nu}}{m}\nabla_{\nu}Z_{\mu}{\rm d}\tau-Z_{\mu}\mathcal{F}_{\text{em}}^{\mu}{\rm d}\tau.$ (37)
Similarly,
${\rm d}\tilde{H}_{\tau}$ $={\rm d}\tilde{E}_{\tau}-\frac{q}{m}\frac{\partial}{\partial x^{\nu}}(Z_{\mu}A^{\mu})\tilde{p}^{\nu}_{\tau}{\rm d}\tau$
$={\rm d}\tilde{E}_{\tau}+Z_{\mu}\mathcal{F}^{\mu}_{\text{em}}{\rm d}\tau-\frac{q}{m}[A_{\mu}\nabla_{\nu}Z^{\mu}+Z^{\mu}\nabla_{\mu}A_{\nu}]\tilde{p}^{\nu}_{\tau}{\rm d}\tau$
$=-Z_{\mu}\left[\xi_{\tau}^{\mu}+\mathcal{F}_{\text{dp}}^{\mu}\right]{\rm d}\tau-\frac{\tilde{p}_{\tau}^{\mu}\tilde{p}_{\tau}^{\nu}}{m}\nabla_{\nu}Z_{\mu}{\rm d}\tau-\frac{q}{m}\pounds_{Z}A_{\mu}\tilde{p}^{\mu}_{\tau}{\rm d}\tau,$ (38)
where £$\pounds_{Z}A_{\mu}$ is the Lie derivative of $A^{\mu}$ along the vector field $Z^{\mu}$. If the electromagnetic field is controlled by an external protocol denoted by $\sigma$, the last term of the above equation can be rewritten as
${\rm d}_{\sigma}\tilde{\mathcal{U}}:=-\frac{q}{m}\pounds_{Z}A_{\mu}\tilde{p}^{\mu}_{\tau}{\rm d}\tau.$ (39)
In the realm of stochastic thermodynamics, the energy exchange between the Brownian particle and the heat reservoir is considered as heat, while the other part of the change of energy of the Brownian particle is considered as work. Therefore, the heat received by the Brownian particle from the heat reservoir is identified to be
${\rm d}\tilde{\mathcal{Q}}_{\tau}:=-Z_{\mu}\left[\xi_{\tau}^{\mu}+\mathcal{F}_{\text{dp}}^{\mu}\right]{\rm d}\tau,$ (40)
and gravitational [26] and electromagnetic works are identified respectively as
${\rm d}\tilde{\mathcal{P}}_{\tau}:=-\frac{\tilde{p}_{\tau}^{\mu}\tilde{p}_{\tau}^{\nu}}{m}\nabla_{\nu}Z_{\mu}{\rm d}\tau,$ (41)
${\rm d}\tilde{\mathcal{W}}_{\tau}:=-Z_{\mu}\mathcal{F}^{\mu}_{\text{em}}{\rm d}\tau.$ (42)
Therefore, the first law of relativistic stochastic thermodynamics can be realized either as
${\rm d}\tilde{E}_{\tau}={\rm d}\tilde{\mathcal{Q}}_{\tau}+{\rm d}\tilde{\mathcal{P}}_{\tau}+{\rm d}\tilde{\mathcal{W}}_{\tau},$ (43)
or as
${\rm d}\tilde{H}_{\tau}={\rm d}\tilde{\mathcal{Q}}_{\tau}+{\rm d}\tilde{\mathcal{P}}_{\tau}+{\rm d}_{\sigma}\tilde{\mathcal{U}}.$ (44)
The energy currents associated to the above two definitions of energy are presented as follows,
$E^{\mu}[\varphi]$ $:=\int_{{(\Gamma^{+}_{m})_{x}}}\eta_{(\Gamma^{+}_{m})_{x}}\frac{p^{\mu}}{m}\varphi E=-Z_{\nu}T^{\mu\nu}[\varphi],$ (45)
$H^{\mu}[\varphi]$ $:=\int_{{(\Gamma^{+}_{m})_{x}}}\eta_{(\Gamma^{+}_{m})_{x}}\frac{p^{\mu}}{m}\varphi H=E^{\mu}[\varphi]-qZ_{\nu}A^{\nu}N^{\mu}[\varphi],$ (46)
wherein the energy-momentum tensor $T^{\mu\nu}[\varphi]$ reads
$T^{\mu\nu}[\varphi]:=\int_{{(\Gamma^{+}_{m})_{x}}}\eta_{(\Gamma^{+}_{m})_{x}}\frac{p^{\mu}p^{\nu}}{m}\varphi.$ (47)
In relativistic kinetic theory, the entropy current associated with classical non-degenerate particles is defined as [28]
$S^{\mu}[\varphi]=-\int_{{(\Gamma^{+}_{m})_{x}}}\eta_{(\Gamma^{+}_{m})_{x}}\frac{p^{\mu}}{m}\varphi\left(\ln\varphi-1\right).$ (48)
In the equilibrium distribution (31), we have the following relation [29]:
$S^{\mu}[\varphi_{\text{eq}}]$ $=\beta^{\mu}P-T^{\mu\nu}[\varphi_{\text{eq}}]\beta^{\nu}+\alpha n^{\mu}[\varphi_{\text{eq}}],$ (49)
where $P:=\frac{1}{d}\Delta_{\mu\nu}(U)T^{\mu\nu}[\varphi_{\text{eq}}]$ is the pressure and $n^{\mu}[\varphi]$ is particle current. Following a standard procedure as did in [29], it can be shown that $T^{\mu\nu}[\varphi_{\text{eq}}]$ and $n^{\mu}[\varphi_{\text{eq}}]$ can both be expressed in terms of the proper velocity $U^{\mu}$ of Bob: $n^{\mu}[\varphi_{\text{eq}}]$ is simply proportional to $U^{\mu}$, while $T^{\mu\nu}[\varphi_{\text{eq}}]$ takes the form of the energy-momentum tensor of a perfect fluid.
Now we would like to reinterpret eq.(49) in terms of the densities of various physical quantities as measured by Alice. This can be achieved by contracting each term in eq.(49) with $Z_{\mu}$. Recalling the fact $\beta^{\mu}=\beta U^{\mu}$, the relationship between the proper velocities of Alice and Bob, can be written as
$U^{\mu}=\gamma_{\rm A}(Z^{\mu}+z^{\mu}),\qquad\gamma_{\rm A}=-U^{\mu}Z_{\mu},\qquad z^{\mu}Z_{\mu}=0,$ (50)
where $\gamma_{\rm A}$ represents the local Lorentz factor arising from the relative motion between Alice and Bob. Consequently, we have
$e_{\rm A}$ $=(\gamma_{\rm A}\beta)^{-1}s_{\rm A}-\alpha(\gamma_{\rm A}\beta)^{-1}n_{\rm A}-P+{\mathcal{T}}_{\rm A},$ (51)
or alternatively
$h_{\rm A}$ $=(\gamma_{\rm A}\beta)^{-1}s_{\rm A}-[qZ_{\mu}A^{\mu}+\alpha(\gamma_{\rm A}\beta)^{-1}]n_{\rm A}-P+{\mathcal{T}}_{\rm A},$ (52)
where
$e_{\rm A}:=T^{\mu\nu}[\varphi_{\text{eq}}]Z_{\mu}Z_{\nu},\quad h_{\rm A}:=e_{\rm A}-qZ_{\mu}A^{\mu}n_{\rm A},\quad n_{\rm A}:=-Z_{\mu}n^{\mu}[\varphi_{\text{eq}}],$ (53)
$s_{\rm A}:=-Z_{\mu}S^{\mu}[\varphi_{\text{eq}}],\quad{\mathcal{T}}_{\rm A}:=-T^{\rho\nu}[\varphi_{\text{eq}}]z_{\mu}Z_{\nu}.$ (54)
$e_{\rm A}$ is the density of the energy $E$, $h_{\rm A}$ is the density of the energy $H$, $n_{\rm A}$ is the particle number density, $s_{\rm A}:=-Z_{\mu}S^{\mu}[\varphi_{\text{eq}}]$ is the entropy density, and ${\mathcal{T}}_{\rm A}$ is the relative kinematic energy density arising from the relative motion between Alice and Bob. The subscript $\rm A$ in the notation for all these density quantities indicates that they are all defined with respect to the observer Alice. Eq. (51) can be viewed as the localized version of the Euler relation in which the coefficients of $s_{\rm A}$ and $n_{\rm A}$ should be interpreted as the temperature and chemical potential as measured by Alice. Therefore we have
$T_{\rm A}:=(\gamma_{\rm A}\beta)^{-1},\qquad\mu_{\rm A}:=-\alpha(\gamma_{\rm A}\beta)^{-1}.$ (55)
Similarly, from eq. (52), we can read off
$\hat{\mu}_{\rm A}:=-[qZ_{\mu}A^{\mu}+\alpha(\gamma_{\rm AB}\beta)^{-1}]=\mu_{\rm A}-qZ_{\mu}A^{\mu},$ (56)
which is recognized to be the electrochemical potential. Inserting eq.(34) into eq. (55), the transformation rule for the temperature and chemical potential presented in Ref.[29] can be easily recovered,
$T_{\rm A}:=(\gamma_{\rm A})^{-1}T_{\rm B},\qquad\mu_{\rm A}:=(\gamma_{\rm A})^{-1}\mu_{\rm B}.$
We can also substitute eqs. (55) and (56) into eqs. (51)-(52) to make the appearance of Euler relation simpler,
$e_{\rm A}$ $=T_{\rm A}s_{\rm A}+\mu_{\rm A}n_{\rm A}-P+{\mathcal{T}}_{\rm A},$ (57)
$h_{\rm A}$ $=T_{\rm A}s_{\rm A}+\hat{\mu}_{\rm A}n_{\rm A}-P+{\mathcal{T}}_{\rm A}.$ (58)
The different choices of definition for the energy induce different chemical potentials, while such choices have no influence on the temperature. Putting these thermodynamic quantities into eq. (31), the equilibrium distribution can be rewritten as
$\varphi_{\text{eq}}={\rm e}^{(E-\mu_{\rm A}-z_{\nu}p^{\nu})/T_{\rm A}}={\rm e}^{(H-\hat{\mu}_{\rm A}-z_{\nu}p^{\nu})/T_{\rm A}}.$ (59)
It remains to introduce the trajectory entropy for the Brownian particle which plays an important role in the formulation of fluctuation theorem. Denoting the phase trajectory of the Brownian particle as $\tilde{Y}_{t}=(\tilde{y}_{t},\tilde{k}_{t})$, the trajectory entropy is defined as
$\tilde{S}_{t}=-\ln\varphi(\tilde{Y}_{t}),$ (60)
The entropy production of the Brownian particle in the time interval $[t_{I},t_{F}]$ is related to the trajectory entropy via
$\Delta S=\int_{\mathcal{S}_{t_{F}}}\eta_{\mathcal{S}_{t_{F}}}Z_{\mu}S^{\mu}[\varphi]-\int_{\mathcal{S}_{t_{I}}}\eta_{\mathcal{S}_{t_{I}}}Z_{\mu}S^{\mu}[\varphi]=\langle\tilde{S}_{t_{F}}\rangle-\langle\tilde{S}_{t_{I}}\rangle.$ (61)
This part of the entropy production is also referred to as the trajectory entropy production.
4 Time-reversal symmetry and its breaking
Before dwelling into the construction of fluctuation theorems in the framework of general relativistic stochastic thermodynamics, it crucial to accurately describe what the TRS is meant on curved spacetime.
In ordinary textbooks on special relativistic field theories, the time-reversal transformation (TRT) is often represented by a coordinate transformation $(x^{0},x^{i})\mapsto(-x^{0},x^{i})$ with different transformation rules for temporal and spatial components of the coordinate bases
$\frac{\partial}{\partial x^{0}}\mapsto-\frac{\partial}{\partial x^{0}},\qquad\frac{\partial}{\partial x^{i}}\mapsto\frac{\partial}{\partial x^{i}}.$ (62)
Since $x^{0}$ serves as the evolution parameter, the spatial components $p^{i}$ of the momentum should reverse their signs, while the temporal component of the momentum remains unchanged, which corresponds to the particle’s energy. Consequently, the transformation under TRT for the contraction of momentum components and the basis obeys
$p=p^{0}\frac{\partial}{\partial x^{0}}+p^{i}\frac{\partial}{\partial x^{i}}\mapsto-p^{0}\frac{\partial}{\partial x^{0}}-p^{i}\frac{\partial}{\partial x^{i}}=-p.$ (63)
To maintain general covariance in our construction, it is better to treat the TRT as a change of the observer’s time orientation, i.e. from the perspective of Alice to that of Carol, rather than treating it as a coordinate transformation. Meanwhile, we adopt the scalar expression like $-Z_{\mu}p^{\mu}$ for defining the energy of a particle instead of using the coordinate dependent definition $p^{0}$. Therefore, we employ an alternative representation of the TRT, which reverses the sign of full momentum (i.e. $p^{\mu}\mapsto-p^{\mu}$) while keeping the spacetime coordinates intact. Such a representation implies the reversal of the particle’s proper time derivative: ${\rm d}/{\rm d}\tau\mapsto-{\rm d}/{\rm d}\tau$. If Alice observes a particle evolving from $\tau_{I}$ to $\tau_{F}$ in its own proper time during the time interval $t_{I}\rightarrow t_{F}$, then Carol will observe the proper time of the charged particle evolves from $\tau_{F}$ to $\tau_{I}$.
Since the momentum is defined only in the local tangent space on curved spacetime, the TRT can be realized as an automorphism of $T_{x}\mathcal{M}$:
$I_{x}:p\mapsto-p,\qquad\text{for }\forall p\in T_{x}\mathcal{M}.$ (64)
Furthermore, $I_{x}$ induces a homeomorphism between the future and past mass shell bundles,
$I:\Gamma^{+}_{m}\to\Gamma^{-}_{m},$ (65)
$I:(x,p)\mapsto(x,I_{x}(p))=(x,-p).$ (66)
Therefore, the phase trajectories lifted by Alice and Carol are on two disconnected regions of the mass shell bundle. For convenience, the trajectory lifted by Alice will be referred to as the forward trajectory, and the one by Carol will be referred to as the reversed trajectory. For later reference, we also list the induced action of the map $I$ on the coordinate basis for the vector field on $T\mathcal{M}$,
$I^{*}\frac{\partial}{\partial p^{\mu}}=-\frac{\partial}{\partial p^{\mu}},\qquad I^{*}\frac{\partial}{\partial x^{\mu}}=\frac{\partial}{\partial x^{\mu}}.$ (67)
Figure 2: The evolution directions of the particle under the perspectives of Alice and Carol.
If we discretize the time interval $[t_{I},t_{F}]$ into a sequence of $N+1$ equal-distance nodes $t_{I}=t_{0}<t_{1}<\cdots<t_{N}=t_{F}$, then the forward trajectory can be written as a sequence $Y_{[t]}:=(Y_{0},Y_{1},\cdots,Y_{N})$ with $Y_{n}:=Y_{t_{n}}$. Consequently, the reversed trajectory can also be written as a sequence $Y_{[t]}^{-}:=(Y_{0}^{-},Y_{1}^{-},\cdots,Y_{N}^{-})$ in which
$Y^{-}_{n}=I(Y_{N-n}).$ (68)
The probability distribution $f=-p^{\mu}Z_{\mu}\varphi/m$ describes the distribution of the intersections of the stochastic trajectories with the space of microstates $\Sigma^{+}_{t}$, and it can be easily extended to the past mass shell bundle by use of the relationship between the forward and the reversed trajectories. Additionally, as a metric-preserving transformation, the TRT $I$ preserves the volume elements on every submanifolds of the tangent bundle $T\mathcal{M}$ mentioned in Sec.  2. Consequently, we have
$I^{*}\varphi=\varphi,\qquad I^{*}\eta_{\Gamma^{+}_{m}}=\eta_{\Gamma^{-}_{m}},\qquad I^{*}\eta_{(\Gamma^{+}_{m})_{x}}=\eta_{(\Gamma^{-}_{m})_{x}},\qquad I^{*}\eta_{\Sigma^{+}_{t}}=\eta_{\Sigma_{t}^{-}},$ (69)
where $\Sigma_{t}^{-}$ is the space of microstates associated to the past directed observer, Carol.
It remains to describe the TRT for the electromagnetic field. Consider the motion of a massive charged particle subjecting to the electromagnetic field $F^{\mu\nu}$. The equation of motion reads
$\frac{{\rm d}p^{\mu}}{{\rm d}\tau}=\frac{q}{m}F^{\mu}{}_{\nu}p^{\nu}.$ (70)
The TRS implies that the above equation of motion is kept invariant. Since $p^{\mu}$ is odd and hence $\frac{{\rm d}p^{\mu}}{{\rm d}\tau}$ is even under the TRT, the invariance of the equation of motion requires either
$q\mapsto q,\qquad F^{\mu\nu}\mapsto-F^{\mu\nu}$ (71)
or
$q\mapsto-q,\qquad F^{\mu\nu}\mapsto F^{\mu\nu}.$ (72)
We will adopt the second convention which is known as Feynman’s convention in the literature[30]. Therefore, we have
$I^{*}q=-q,\qquad I^{*}\mathcal{F}_{\text{em}}^{\mu}=\mathcal{F}_{\text{em}}^{\mu}.$ (73)
Using the definition (26), it is not difficult to check that the induced action of the TRT on the Liouville vector field is given by
$I^{*}\mathscr{L}_{F}=-\mathscr{L}_{F}.$ (74)
Bringing all the above conventions together, we will see that the equations for the phase trajectory of a massive charged particle moving in a generic curved spacetime and subjecting to an external electromagnetic field is kept invariant under the TRT,
$\frac{{\rm d}x^{\mu}}{{\rm d}\tau}$ $=\frac{p^{\mu}}{m},$ (75)
$\frac{{\rm d}p^{\mu}}{{\rm d}\tau}$ $=\frac{q}{m}F^{\mu}{}_{\nu}p^{\nu}-\frac{1}{m}\varGamma^{\mu}{}_{\alpha\beta}p^{\alpha}p^{\beta}.$ (76)
Let us now bring the above system into a broader picture. Consider a scenario in which there are a great number of massive charged particles of different masses and charges moving together. The electromagnetic field is produced by the particles themselves and the spacetime geometry is determined by the masses and charges of the particles. We assume that the system is consisted of different species of particles, each species carries different mass $m_{s}$ and charge $q_{s}$ and obey a different distribution $\Phi_{s}$, which are differentiated from each other by the suffix $s$. We also assume that the distributions $\Phi_{s}$ are invariant under the TRT, i.e. $I^{*}\Phi_{s}=\Phi_{s}$. Then the total electric current
$J^{\mu}=\sum_{s}\int\eta_{(\Gamma_{m_{s}}^{+})_{x}}\frac{q_{s}p^{\mu}_{s}}{m_{s}}\Phi_{s}$ (77)
as well as the energy-momentum tensor contributed by the particles
$T^{\mu\nu}_{\text{pa}}=\sum_{s}\int\eta_{(\Gamma_{m_{s}}^{+})_{x}}\frac{p^{\mu}_{s}p^{\nu}_{s}}{m_{s}}\Phi_{s}$ (78)
should be both TRT invariants. Consequently, the Maxwell equation
$\nabla_{\nu}F^{\mu\nu}=J^{\mu}$ (79)
that determines the electromagnetic field as well as the Einstein equation
$R^{\mu\nu}-\frac{1}{2}g^{\mu\nu}R=8\pi G\left(T^{\mu\nu}_{\text{pa}}+T^{\mu\nu}_{\text{em}}\right)$ (80)
that determines the spacetime geometry should all be invariant under the TRT, wherein
$T^{\mu\nu}_{\text{em}}=F^{\mu\rho}F^{\nu}{}_{\rho}-\frac{1}{4}g^{\mu\nu}F^{\rho\sigma}F_{\rho\sigma}$ (81)
is the energy-momentum tensor of the electromagnetic field, whose TRT invariance is self-evident.
Stochastic mechanics can be regarded as an effective theory for describing a complete mechanical system within specific spatial and temporal scales [31, 32, 33]. Let us consider a heavy particle in the above system, assuming that the remaining particles constitute a heat reservoir which has already reached equilibrium. The electromagnetic interaction acting on the heavy particle can be divided into two parts: the coarse-grained averaging effects at larger spatial and temporal scales, and the stochastic remnants at smaller scales. Consequently, eqs. (75)-(76) can be approximated by LEτ [1], and this heavy particle is the Brownian particle.
However, such a coarse-grained description violates the TRS. Let us first consider the damping force. If the forward and reversed trajectories simultaneously satisfy following equation
$\frac{1}{m}p^{\nu}\nabla_{\nu}p^{\mu}=\mathcal{K}^{\mu\nu}U_{\nu},$ (82)
the damping tensor must reverse its sign under the TRT, since the velocity of the heat reservoir should reverse its sign. The sign change in $\mathcal{K}^{\mu\nu}$ implies that the damping force reverses its role from a decelerating force to an accelerating force, which makes a difference between future- and past-directed observers. On the other hand, the Einstein relation [2]
$\mathcal{D}^{\mu\nu}=2T_{\rm B}\mathcal{K}^{\mu\nu}$ (83)
requires that the damping tensor should be kept invariant under the TRT, because the diffusion tensor $\mathcal{D}^{\mu\nu}=\mathcal{R}^{\mu}{}_{\mathfrak{a}}\mathcal{R}^{\nu}{}_{\mathfrak{a}}$ is a quadratic form in $\mathcal{R}^{\mu}{}_{\mathfrak{a}}$ whose eigenvalues must be non-negative. Therefore, the assumption that eq.(82) is invariant under the TRT has to be wrong. The correct behaviors of the damping and diffusion tensors and the stochastic amplitudes are provided as follows,
$I^{*}\mathcal{K}^{\mu\nu}=\mathcal{K}^{\mu\nu},\qquad I^{*}\mathcal{D}^{\mu\nu}=\mathcal{D}^{\mu\nu},\qquad I^{*}\mathcal{R}^{\mu}{}_{\mathfrak{a}}=\pm\mathcal{R}^{\mu}{}_{\mathfrak{a}}.$ (84)
The sign of the stochastic amplitude $\mathcal{R}^{\mu}{}_{\mathfrak{a}}$ is of little significance, since the probability distribution of the Gaussian noise is an even function. For convenience, we adopt $I^{*}\mathcal{R}^{\mu}{}_{\mathfrak{a}}=\mathcal{R}^{\mu}{}_{\mathfrak{a}}$. For the sake of mathematical consistency, we define a vector field $V^{a}$ on $T\mathcal{M}$, which satisfies
$V^{a}|_{(x,p)}=\begin{cases}\displaystyle U^{\mu}|_{x}\frac{\partial}{\partial p^{\mu}}\bigg{|}_{(x,p)}&(x,p)\in\Gamma^{+}_{m}\\ \displaystyle-U^{\mu}|_{x}\frac{\partial}{\partial p^{\mu}}\bigg{|}_{(x,p)}&(x,p)\in\Gamma^{-}_{m}\end{cases}.$ (85)
Then the damping force and the right-hand side of eq. (27) can be rewritten as
$\mathcal{F}_{\text{dp}}^{a}=\mathcal{K}^{ab}V_{b},\qquad\mathcal{I}^{a}[\varphi]=\frac{1}{2}\mathcal{D}^{ab}\nabla^{(h)}_{b}\varphi-\mathcal{K}^{ab}V_{b}\varphi,$ (86)
and it is easy to check that their transformation rules under the TRT are
$I^{*}\mathcal{F}_{\text{dp}}=\mathcal{F}_{\text{dp}},\qquad I^{*}\mathcal{I}[\varphi]=\mathcal{I}[\varphi].$ (87)
Therefore, the probability current (29) of the Brownian particle can be divided into the even and odd parts under the TRT, i.e.
$\mathscr{J}_{\text{r}}[\varphi]=\dfrac{\varphi}{m}\mathscr{L}_{F},\qquad\mathscr{J}_{\text{d}}[\varphi]=-\mathcal{I}[\varphi],$ (88)
with
$I^{*}\mathscr{J}_{\text{r}}[\varphi]=-\mathscr{J}_{\text{r}}[\varphi],\qquad I^{*}\mathscr{J}_{\text{d}}[\varphi]=\mathscr{J}_{\text{d}}[\varphi].$ (89)
It is evident that $\mathscr{J}_{\text{d}}[\varphi]$ violates the TRS. The entropy production is always closely related to the breaking of the TRS, and we have proved [3] that, on the ensemble level,
$\nabla_{\mu}S^{\mu}[\varphi]=-\int\eta_{(\Gamma_{m}^{+})_{x}}\varphi^{-1}\frac{\partial\varphi}{\partial p^{\mu}}\mathscr{J}^{\mu}_{\mathrm{d}}[\varphi],$ (90)
$\nabla_{\mu}S_{R}^{\mu}=\int\eta_{(\Gamma_{m}^{+})_{x}}\beta_{\mu}\mathscr{J}^{\mu}_{\text{d}}[\varphi],$ (91)
where $S^{\mu}[\varphi]$ and $S_{R}^{\mu}$ respectively denote the entropy currents of the Brownian particle and of the heat reservoir. Eq. (91) is actually the relativistic version of the celebrated Clausius’ identity
$\nabla_{\mu}S^{\mu}_{R}=-\frac{Q_{\rm B}[\varphi]}{T_{\rm B}}=\frac{Q_{R}}{T_{\rm B}},$ (92)
where $Q_{\rm B}[\varphi]$ is the heat transfer rate from Bob’s perspective, and $Q_{R}$ is that of the heat reservoir. It is worth noticing that the Clausius’ identity holds only from the perspective of Bob.
5 Fluctuation theorem
Unlike the case of ordinary differential equations, there is no deterministic solution for stochastic differential equation. The best one can do is to determine the probability for a certain trajectory to be realized. Consequently, the breaking of the TRS of the Langevin equation could be described in terms of the non-equal probabilities for the forward and reversed trajectories,
$\Pr[\tilde{Y}_{[t]}=Y_{[t]}]\neq\Pr[\tilde{Y}_{[t]}^{-}=Y_{[t]}^{-}],$ (93)
where $\tilde{Y}_{[t]}=(\tilde{Y}_{0},\tilde{Y}_{1},\cdots,\tilde{Y}_{N})$ and $\tilde{Y}_{[t]}^{-}=(\tilde{Y}^{-}_{0},\tilde{Y}_{1}^{-},\cdots,\tilde{Y}_{N}^{-})$ denote the forward and reversed processes, and $Y_{[t]}$ and $Y^{-}_{[t]}$ denote the forward and reversed trajectories which are already described in Sec. 4. Notice that “forward process” and “forward trajectory” are different concepts: the latter is a concrete realization of the former. Unlike the relation between forward and reversed trajectories, the only requirement on the reversed process is that its initial state is identical to the TRT of the final state of the forward process, i.e.
$\tilde{Y}^{-}_{0}=I(\tilde{Y}_{N}).$ (94)
The Lorentz factors relative to Alice and Carol can be related via the TRT
$\gamma_{\rm A}=-\lambda Z_{\mu}p^{\mu}/m=\lambda C_{\mu}p^{\mu}/m=I^{*}\gamma_{C}.$ (95)
These Lorentz factors can be merged into a single scalar field on the complete mass shell bundle,
$\gamma:=\begin{cases}\gamma_{\rm A}|_{Y}&Y\in\Gamma^{+}_{m}\\ \gamma_{C}|_{Y}&Y\in\Gamma^{-}_{m}\end{cases},$ (96)
which is even under the TRT, $I^{*}\gamma=\gamma$. Accounting for the transformation rules under the TRT, we can rearrange the spatial components of the LEt presented in eqs. (22)-(23) into the form
${\rm d}\tilde{y}_{t}^{i}$ $=\frac{\tilde{k}^{i}_{t}}{m}\gamma^{-1}{\rm d}t,$ (97)
${\rm d}\tilde{k}_{t}^{i}$ =^$=\hat{\mathcal{R}}^{i}{}_{\mathfrak{a}}\circ_{S}{\rm d}\tilde{W}^{\mathfrak{a}}_{n}+F^{i}{\rm d}t+\bar{F}^{i}{\rm d}t,$ (98)
where ^$\hat{\mathcal{R}}^{i}{}_{\mathfrak{a}}:=\gamma^{-1/2}\mathcal{R}^{i}{}_{\mathfrak{a}}$ and
$F^{i}:=\gamma^{-1}\left(\mathcal{F}^{i}_{\text{em}}-\frac{1}{m}\varGamma^{i}{}_{\alpha\beta}\tilde{k}^{\alpha}\tilde{k}^{\beta}\right),\quad\bar{F}^{i}$ $:=\gamma^{-1}\left(\mathcal{F}_{\text{add}}^{i}-\frac{1}{2}\mathcal{D}^{ij}\gamma^{1/2}\nabla^{(h)}_{j}\gamma^{-1/2}+\mathcal{F}^{i}_{\text{dp}}\right),$ (99)
which have opposite behaviors under the TRT,
$F^{i}|_{Y}=F^{i}|_{I(Y)},\qquad\bar{F}^{i}|_{Y}=-\bar{F}^{i}|_{I(Y)}.$ (100)
In order to describe the fluctuation theorems at the trajectory level, we need a discretized version of LEt, which reads
${\rm d}\tilde{y}_{n}^{i}$ $=\frac{\tilde{k}^{i}_{n}}{m}\gamma^{-1}|_{\tilde{Y}_{\bar{n}}}{\rm d}t,$ (101)
${\rm d}\tilde{k}_{n}^{i}$ =^$=\hat{\mathcal{R}}^{i}{}_{\mathfrak{a}}|_{\tilde{Y}_{\bar{n}}}{\rm d}\tilde{W}^{\mathfrak{a}}_{n}+F^{i}|_{\tilde{Y}_{\bar{n}}}{\rm d}t+\bar{F}^{i}|_{\tilde{Y}_{\bar{n}}}{\rm d}t,$ (102)
where $\tilde{Y}_{\bar{n}}:=(\tilde{Y}_{n+1}+\tilde{Y}_{n})/2$ which comes from the Stratonovich coupling. We also need to introduce the measures on the space of forward and reversed trajectories which is defined to be the successive wedge products of the volume element of the space of microstates at each moment of the discrete time,
$\mathcal{D}[Y_{[t]}]:=\eta_{\Sigma_{0}}\wedge\eta_{\Sigma_{1}}\wedge\cdots\wedge\eta_{\Sigma_{N}},\qquad\mathcal{D}[Y^{-}_{[t]}]:=\eta_{\Sigma^{-}_{0}}\wedge\eta_{\Sigma^{-}_{1}}\wedge\cdots\wedge\eta_{\Sigma^{-}_{N}}.$ (103)
Eq. (69) implies that these measures are images of each other under the TRT,
$I^{*}\mathcal{D}[Y_{[t]}]=\mathcal{D}[Y^{-}_{[t]}].$ (104)
In other words, $Y_{[t]}\mapsto Y^{-}_{[t]}$ is a volume preserving map.
Since both the forward and reversed processes are Markovian, the trajectory probabilities can be written as products of transition probabilities with the initial probability,
$\Pr[\tilde{Y}_{[t]}=Y_{[t]}]$ $=\bigg{(}\prod_{n=0}^{N-1}\Pr[\tilde{Y}_{n+1}=Y_{n+1}|\tilde{Y}_{n}=Y_{n}]\bigg{)}\Pr[\tilde{Y}_{0}=Y_{0}],$ (105)
$\Pr[\tilde{Y}^{-}_{[\bar{t}\,]}=Y^{-}_{[t]}]$ $=\bigg{(}\prod_{n=0}^{N-1}\Pr[\tilde{Y}^{-}_{n+1}=Y^{-}_{n+1}|\tilde{Y}^{-}_{n}=Y^{-}_{n}]\bigg{)}\Pr[\tilde{Y}^{-}_{0}=Y^{-}_{0}]$
$=\bigg{(}\prod_{n=0}^{N-1}\Pr[\tilde{Y}^{-}_{n+1}=I(Y_{N-n-1})|\tilde{Y}^{-}_{n}=I(Y_{N-n})]\bigg{)}\Pr[I(\tilde{Y}_{N})=I(Y_{N})]$
$=\bigg{(}\prod_{n=0}^{N-1}\Pr[\tilde{Y}^{-}_{N-n}=I(Y_{n})|\tilde{Y}^{-}_{N-n-1}=I(Y_{n+1})]\bigg{)}\Pr[\tilde{Y}_{N}=Y_{N}].$ (106)
Eq. (68) and eq. (94) are used in the second line of eq. (106), and the volume preserving property of $I$ is used in the third line. Appendix B proves the following continuum limit
$\lim_{N\rightarrow+\infty}\prod_{n=0}^{N-1}\frac{\Pr[\tilde{Y}_{n+1}=Y_{n+1}|\tilde{Y}_{n}=Y_{n}]}{\Pr[\tilde{Y}^{-}_{N-n}=I(Y_{n})|\tilde{Y}^{-}_{N-n-1}=I(Y_{n+1})]}$
$\qquad=\frac{(\lambda\gamma^{-1})|_{Y_{0}}}{(\lambda\gamma^{-1})|_{Y_{N}}}\exp\left[\int^{t_{F}}_{t_{I}}\gamma^{-1}{\rm d}t\frac{1}{T_{\rm B}}({m}a^{\mu}-\mathcal{F}^{\mu}_{\text{em}})U_{\mu}\right].$ (107)
Under the perspective of Bob, the complete differential of the energy is
${\rm d}E_{\rm B}=-\frac{p^{\nu}}{m}\nabla_{\nu}(p^{\mu}U_{\mu}){\rm d}\tau=-ma^{\mu}U_{\mu}{\rm d}\tau+{\rm d}\mathcal{P},$ (108)
or
${\rm d}H_{\rm B}=-\frac{p^{\nu}}{m}\nabla_{\nu}(P^{\mu}U_{\mu}){\rm d}\tau=-ma^{\mu}U_{\mu}{\rm d}\tau+{\rm d}\mathcal{P}+{{\rm d}_{\sigma}\mathcal{U}}-{\rm d}\mathcal{W}.$ (109)
Comparing with eq. (43), it is clear that the integral on the exponent in eq. (5) is actually the increase of the entropy of the heat reservoir
$\int\gamma^{-1}{\rm d}t\frac{1}{T_{\rm B}}(ma^{\mu}-\mathcal{F}^{\mu}_{\text{em}})U_{\mu}=-\int\frac{{\rm d}\mathcal{Q}}{T_{\rm B}}=\Delta S_{R},$ (110)
where the second equality encodes the Clausius equality at the level of trajectories. Let us stress that, although the Clausius equality holds only from the perspective of Bob, the total increase of the entropy of the heat reservoir is actually observer-independent. This fact enables us to make use of eq. (110) in addressing the fluctuation theorem from the perspective of an arbitrary observer.
In the continuum limit, the logarithm of the ratio of the probabilities of the forward and reversed trajectories is the sum of the trajectory entropy production of the Brownian particle with the entropy increase of the heat reservoir,
$\ln\frac{\Pr[\tilde{Y}_{[t]}=Y_{[t]}]}{\Pr[\tilde{Y}^{-}_{[t]}=Y^{-}_{[t]}]}$ $=\lim_{N\to\infty}\sum_{n=0}^{N-1}\ln\frac{\Pr[\tilde{Y}_{n+1}=Y_{n+1}|\tilde{Y}_{n}=Y_{n}]}{\Pr[\tilde{Y}^{-}_{N-n}=I(Y_{n})|\tilde{Y}^{-}_{N-n-1}=I(Y_{n+1})]}$
$\qquad+\ln\frac{(\lambda\gamma^{-1})|_{X_{0}}\Pr[\tilde{Y}_{0}=Y_{0}]}{(\lambda\gamma^{-1})|_{Y_{N}}\Pr[\tilde{Y}_{N}=Y_{N}]}$
$=-\int\frac{{\rm d}\mathcal{Q}}{T_{\rm B}}+\ln\frac{\varphi(Y_{0})}{\varphi(Y_{N})}$
$=\Delta S_{R}+\Delta S.$ (111)
Please be reminded that the relation eq. (30) between different distribution functions and the definition (60) of the trajectory entropy production are used here. Denoting the total entropy production at the trajectory level as $\Sigma_{Y_{[t]}}:=\Delta S_{R}+\Delta S$, the usual form of the detailed fluctuation theorem arises
$\frac{\Pr[\tilde{Y}_{[t]}=Y_{[t]}]}{\Pr[\tilde{Y}^{-}_{[t]}=Y^{-}_{[t]}]}={\rm e}^{\Sigma_{Y_{[t]}}}.$ (112)
This theorem reveals how the breaking of the TRS is associated with the entropy production at the trajectory level. Moreover, taking the expectation value of ${\rm e}^{\Sigma_{X_{[t]}}}$ and using Jensen inequality, one gets the integral fluctuation theorem
${\rm e}^{-\left\langle\Sigma_{\tilde{Y}_{[t]}}\right\rangle}\leq\left\langle{\rm e}^{-\Sigma_{\tilde{Y}_{[t]}}}\right\rangle=\int\mathcal{D}[Y_{[t]}]\Pr[\tilde{Y}^{-}_{[t]}=Y^{-}_{[t]}]=\int\mathcal{D}[Y^{-}_{[t]}]\Pr[\tilde{Y}^{-}_{[t]}=Y^{-}_{[t]}]=1,$ (113)
which implies that the statistical expectation value of the total entropy production must be no-negative, i.e. $\left\langle\Sigma_{\tilde{Y}_{[t]}}\right\rangle\geq 0$.
6 Concluding remarks
Fluctuation theorems not only are important in understanding the theoretical origin of irreversibility, but also have found various applications in diverse areas ranging from macroscopic to mesoscopic systems, and even to active matter. The research carried our in the present work opens new area for the potential applications of fluctuation theorems which embodies relativistic gravity, as in typical cases of cosmological and astrophysical processes.
In our construction, it is important to decouple the choice of observers from the choice of coordinate systems, which has several important consequences. First, it allows us to parametrize the stochastic trajectory using the proper time of the observer field, which helps to get rid of the random clock carried by the Brownian particle. Second, it helps to properly describe the TRT in a coordinate independent manner. Last, it is precisely such decoupling that makes the construction fully general covariant.
Although the values of many thermodynamic quantities are observer dependent, including e.g. the energy and its density, the temperature and chemical potential, and even the Clausius’ identity etc., the total entropy production is not among the observer dependent quantities. Meanwhile, the trajectory probability is a pure mathematical entity which is also observer independent. Therefore, it is not surprising that the form of the fluctuation theorems obtained in this work is identical to that obtained from non-relativistic stochastic thermodynamics.
Acknowledhement
This work is supported by the National Natural Science Foundation of China under the grant No. 12275138.
Conflict of interests declaration
The authors declare no known conflict of interests.
Appendix A Coordinate bases for tangent vectors
In our description of the tangent bundle $T\mathcal{M}$, we employ $22d+$ 2 independent coordinates $(x^{\mu},p^{\mu})$ with the corresponding coordinate basis $\left(\left.\frac{\partial}{\partial x^{\mu}}\right|_{T\mathcal{M}},\left.\frac{\partial}{\partial p^{\mu}}\right|_{T\mathcal{M}}\right)$ for tangent vectors on $T\mathcal{M}$. However, due to the mass shell constraint, the above coordinates are no longer independent on the mass shell bundle $\Gamma^{+}_{m}$, rendering the coordinate basis also redundant. The same also happens when we restrict ourselves from the mass shell bundle to the space of microstates $\Sigma_{t}^{+}$. The aim of this appendix is to resolve the redundancies in the coordinate bases on various useful submanifolds of $T\mathcal{M}$.
To resolve the coordinate redundancy on $\Gamma^{+}_{m}$, we take the coordinates on $\Gamma^{+}_{m}$ to be $(x^{\mu},p^{i})$ and view $p^{0}$ as a function
$p^{0}(x^{\mu},p^{i})=\frac{g_{0i}(x)p^{i}\pm\sqrt{\left[g_{0i}(x)p^{i}\right]^{2}-g_{00}(x)\left[m^{2}+g_{ij}(x)p^{i}p^{j}\right]}}{-g_{00}(x)}.$ (114)
This leads to the partial derivatives
$\frac{\partial p^{0}}{\partial x^{\mu}}=-\frac{1}{2p_{0}}\frac{\partial g_{\alpha\beta}}{\partial x^{\mu}}p^{\alpha}p^{\beta},\qquad\frac{\partial p^{0}}{\partial p^{i}}=-\frac{p_{i}}{p_{0}}.$ (115)
It is essential to recognize that the coordinate basis for tangent vectors on $\Gamma^{+}_{m}$ is different from the corresponding subset of coordinate basis on $T\mathcal{M}$,
$\left.\frac{\partial}{\partial x^{\mu}}\right|_{T\mathcal{M}}\neq\left.\frac{\partial}{\partial x^{\mu}}\right|_{\Gamma^{+}_{m}},\quad\left.\frac{\partial}{\partial p^{i}}\right|_{T\mathcal{M}}\neq\left.\frac{\partial}{\partial p^{i}}\right|_{\Gamma^{+}_{m}}.$ (116)
To demonstrate this difference, let us consider a scalar field $f(x^{\mu},p^{\mu})$ on $T\mathcal{M}$ and its restriction on $\Gamma^{+}_{m}$, i.e.
$\left.f\right|_{\Gamma^{+}_{m}}(x^{\mu},p^{i}):=f(x^{\mu},p^{0}(x^{\mu},p^{i}),p^{i}).$ (117)
The action of the coordinate basis vectors on $\left.f\right|_{\Gamma^{+}_{m}}$ reads
$\left.\frac{\partial}{\partial p^{i}}\right|_{\Gamma^{+}_{m}}\left.f\right|_{\Gamma^{+}_{m}}$ $=\left.\frac{\partial}{\partial p^{i}}\right|_{T\mathcal{M}}f+\frac{\partial p^{0}}{\partial p^{i}}\left.\frac{\partial}{\partial p^{0}}\right|_{T\mathcal{M}}f=\left.\frac{\partial}{\partial\breve{p}^{i}}\right|_{T\mathcal{M}}f,$ (118)
$\left.\frac{\partial}{\partial x^{\mu}}\right|_{\Gamma^{+}_{m}}\left.f\right|_{\Gamma^{+}_{m}}$ $=\left.\frac{\partial}{\partial x^{\mu}}\right|_{T\mathcal{M}}f+\frac{\partial p^{0}}{\partial x^{\mu}}\left.\frac{\partial}{\partial p^{0}}\right|_{T\mathcal{M}}f=\left.\frac{\partial}{\partial\breve{x}^{\mu}}\right|_{T\mathcal{M}}f,$ (119)
where the “breved” partial derivatives are defined as
$\left.\frac{\partial}{\partial\breve{p}^{i}}\right|_{T\mathcal{M}}:=\left[\frac{\partial}{\partial p^{i}}-\frac{p_{i}}{p_{0}}\frac{\partial}{\partial p^{0}}\right]_{T\mathcal{M}},\qquad\left.\frac{\partial}{\partial\breve{x}^{\mu}}\right|_{T\mathcal{M}}:=\left[e_{\mu}+\varGamma^{i}{}_{\beta\mu}p^{\beta}\frac{\partial}{\partial\breve{p}^{i}}\right]_{T\mathcal{M}}.$ (120)
Therefore, the partial derivatives on $\Gamma^{+}_{m}$ and $T\mathcal{M}$ are related via
$i^{*}\left.\frac{\partial}{\partial p^{i}}\right|_{\Gamma^{+}_{m}}=\left.\frac{\partial}{\partial\breve{p}^{i}}\right|_{T\mathcal{M}},\qquad i^{*}\left.\frac{\partial}{\partial x^{\mu}}\right|_{\Gamma^{+}_{m}}=\left.\frac{\partial}{\partial\breve{x}^{\mu}}\right|_{T\mathcal{M}},$ (121)
where $i:\Gamma^{+}_{m}\rightarrow T\mathcal{M}$ is the embedding map. From the exterior geometric point of view, we can safely identify $\{\partial/\partial\breve{x}^{\mu},\partial/\partial\breve{p}^{i}\}$ as the coordinate basis on the mass shell bundle, omitting the subscripts.
The tangent vector $\mathcal{V}=\mathcal{V}^{\mu}(\partial/{\partial p^{\mu}})$ of the momentum space must be orthogonal to the normal vector $\hat{N}=p^{\mu}(\partial/\partial p^{\mu})/m$. The orthogonality condition $\mathcal{V}^{\mu}p_{\mu}=0$ implies that
$\mathcal{V}^{\mu}\frac{\partial}{\partial p^{\mu}}=\mathcal{V}^{i}\frac{\partial}{\partial p^{i}}-\frac{\mathcal{V}^{i}p_{i}}{p_{0}}\frac{\partial}{\partial p^{0}}=\mathcal{V}^{i}\frac{\partial}{\partial\breve{p}^{i}}.$ (122)
Therefore, the tangent vectors of the momentum space have two different representations under the above two bases. We will use them interchangeably. There is a similar property for the vectors on mass shell bundle. Let $\mathscr{V}$ be a tangent vector of the phase trajectory $X_{\tau}:=(x_{\tau},p_{\tau})$. As a vector on tangent bundle, $\mathscr{V}$ can be expanded as
$\mathscr{V}=\frac{p^{\mu}}{m}\frac{\partial}{\partial x^{\mu}}+m\mathcal{A}^{\mu}\frac{\partial}{\partial p^{\mu}}=\frac{p^{\mu}}{m}e_{\mu}+ma^{\mu}\frac{\partial}{\partial p^{\mu}},$ (123)
where $\mathcal{A}^{\mu}:=m^{-1}{\rm d}p^{\mu}/{\rm d}\tau$ is the coordinate acceleration and $a^{\mu}:=\mathcal{A}^{\mu}+\varGamma^{\mu}{}_{\alpha\beta}p^{\alpha}p^{\beta}/m^{2}$ is the covariant acceleration. The mass shell condition implies that the covariant acceleration must always be orthogonal to the momentum, so that $\mathscr{V}$ can be rewritten as
$\mathscr{V}=\frac{p^{\mu}}{m}e_{\mu}+ma^{i}\frac{\partial}{\partial\breve{p}^{i}}=\frac{p^{\mu}}{m}\frac{\partial}{\partial\breve{x}^{\mu}}+m\mathcal{A}^{i}\frac{\partial}{\partial\breve{p}^{i}}.$ (124)
The two representations for vectors on the mass shell bundle under the bases $\{\partial/\partial\breve{x}^{\mu},\partial/\partial\breve{p}^{i}\}$ and $\{\partial/\partial x^{\mu},\partial/\partial p^{\mu}\}$ will also be used interchangeably.
The space of microstates $\Sigma^{+}_{t}$ is also an embedding submanifold of the mass shell bundle defined by fixing $t(x)$ to be a constant. Therefore, the coordinates on $\Gamma^{+}_{m}$ becomes redundant once again on $\Sigma^{+}_{t}$. Such redundancy can also be eliminated by taking $(x^{i},p^{i})$ as coordinate of $\Sigma^{+}_{t}$ and viewing $x^{0}(t,x^{i})$ as a function. The restriction of $f$ on $\Sigma^{+}_{t}$ is defined as
$\left.f\right|_{\Sigma_{t}^{+}}(x^{i},p^{i}):=\left.f\right|_{\Gamma^{+}_{m}}(x^{0}(t,x^{i}),x^{i},p^{i})=f(x^{0}(t,x^{i}),x^{i},p^{0}(x^{0}(t,x^{i}),x^{i},p^{i}),p^{i}).$ (125)
Correspondingly, the partial derivatives acting on $f|_{\Sigma^{+}_{t}}$ can be evaluated to be
$\left.\frac{\partial}{\partial x^{i}}\right|_{\Sigma_{t}^{+}}\left.f\right|_{\Sigma_{t}^{+}}$ $=\left.\frac{\partial}{\partial x^{i}}\right|_{\Gamma^{+}_{m}}\left.f\right|_{\Gamma^{+}_{m}}+\frac{\partial x^{0}}{\partial x^{i}}\left.\frac{\partial}{\partial x^{0}}\right|_{\Gamma^{+}_{m}}\left.f\right|_{\Gamma^{+}_{m}}=\left.\frac{\partial}{\partial\breve{x}^{i}}\right|_{T\mathcal{M}}f-\frac{\partial_{i}t}{\partial_{0}t}\left.\frac{\partial}{\partial\breve{x}^{0}}\right|_{T\mathcal{M}}f,$ (126)
$\left.\frac{\partial}{\partial p^{i}}\right|_{\Sigma_{t}^{+}}\left.f\right|_{\Sigma^{+}_{t}}$ $=\left.\frac{\partial}{\partial p^{i}}\right|_{\Gamma^{+}_{m}}\left.f\right|_{\Gamma^{+}_{m}}=\left.\frac{\partial}{\partial\breve{p}^{i}}\right|_{T\mathcal{M}}f,$ (127)
$\frac{\partial}{\partial t}\left.f\right|_{\Sigma^{+}_{t}}$ $=\frac{\partial x^{0}}{\partial t}\left.\frac{\partial}{\partial x^{0}}\right|_{\Gamma^{+}_{m}}\left.f\right|_{\Gamma^{+}_{m}}=\frac{1}{\partial_{0}t}\left.\frac{\partial}{\partial\breve{x}^{0}}\right|_{T\mathcal{M}}f,$ (128)
thanks to the relations (118)-(119). Let $\pi_{t}:\Sigma_{t}^{+}\rightarrow\Gamma^{+}_{m}$ be the embedding map and define
$\left.\frac{\partial}{\partial\hat{x}^{i}}\right|_{T\mathcal{M}}:=\left.\frac{\partial}{\partial\breve{x}^{i}}\right|_{T\mathcal{M}}-\frac{\partial_{i}t}{\partial_{0}t}\left.\frac{\partial}{\partial\breve{x}^{0}}\right|_{T\mathcal{M}},$ (129)
we have
$(i\circ\pi_{t})^{*}\left.\frac{\partial}{\partial p^{i}}\right|_{\Sigma_{t}^{+}}=\left.\frac{\partial}{\partial\breve{p}^{i}}\right|_{T\mathcal{M}},\qquad(i\circ\pi_{t})^{*}\left.\frac{\partial}{\partial x^{i}}\right|_{\Sigma_{t}^{+}}=\left.\frac{\partial}{\partial\hat{x}^{i}}\right|_{T\mathcal{M}}.$ (130)
Therefore, we can take $\{\partial/\partial\hat{x}^{i},\partial/\partial\breve{p}^{i}\}$ as the coordinate basis on $\Sigma^{+}_{t}$ without specifying the manifold with a subscript.
Since the phase trajectory does not lie on a certain $\Sigma^{+}_{t}$, the basis $\{\partial/\partial\hat{x}^{i},\partial/\partial\breve{p}^{i}\}$ is insufficient to describe the tangent vector of the phase trajectory. The missing dimension in the tangent vector of the phase trajectory is described by $\partial/\partial t$ as described in eq. (128), which, from the exterior geometric point of view, can be simply denoted
$\frac{\partial}{\partial t}=\frac{1}{\partial_{0}t}\frac{\partial}{\partial\breve{x}^{0}}.$ (131)
Therefore, under the above vector basis, the tangent vector (124) to the phase trajectory can be re-expressed as
$\mathscr{V}=\gamma\frac{\partial}{\partial t}+\frac{p^{i}}{m}\frac{\partial}{\partial\hat{x}^{i}}+m\mathcal{A}^{i}\frac{\partial}{\partial\breve{p}^{i}},$ (132)
where
$\gamma=\frac{{\rm d}t}{{\rm d}\tau}=\frac{1}{m}p^{\mu}\partial_{\mu}t$ (133)
is the local Lorentz factor between the particle and the observer Alice.
Appendix B Continuum limit
This appendix provides the details for proving eq. (5).
First, we introduce two mathematical lemmas.
1. Let $A$ be a full rank square matrix and $B$ be an arbitrary matrix of the same size. Then the determinant of $A+B{\rm d}t+o(t^{2})$ can be expanded into power series in ${\rm d}t$,
$\det[A+B{\rm d}t+o({\rm d}t^{2})]=\det[A]+\det[A]\mathrm{Tr}[A^{-1}B]{\rm d}t+o({\rm d}t^{2}).$ (134)
2. Let $f(t)$ be a continuous function on $[t_{I},t_{F}]$, then
$\lim_{N\rightarrow+\infty}\prod_{n=0}^{N-1}[1+f(t_{n}){\rm d}t+o({\rm d}t^{2})]=\exp\left[\int_{t_{I}}^{t_{F}}f(t){\rm d}t\right],$ (135)
where ${\rm d}t=(t_{F}-t_{I})/N$ and $t_{n}\in[t_{I}+n{\rm d}t,t_{I}+(n+1){\rm d}t]$.
Combining the above two lemmas, the following corollary can be deduced,
$\lim_{N\rightarrow+\infty}\prod_{n=0}^{N-1}\frac{\det[A(t_{n})+B(t_{n}){\rm d}t+o({\rm d}t^{2})]}{\det[A(t_{n})+C(t_{n}){\rm d}t+o({\rm d}t^{2})]}=\exp\left[\int_{t_{I}}^{t_{F}}\mathrm{Tr}[A^{-1}(B-C)]{\rm d}t\right],$ (136)
where $A(t)$, $B(t)$ and $C(t)$ are all time-dependent matrix functions.
Let $\mathscr{O}$ be a function on the mass shell bundle and $Y_{t}=X_{\tau(t)}$ be a phase trajectory parametrized by the proper time of Alice. Then
$\frac{\mathscr{O}(Y_{t_{F}})}{\mathscr{O}(Y_{t_{I}})}$ $=\frac{\mathscr{O}(X_{\tau_{F}})}{\mathscr{O}(X_{\tau_{I}})}=\lim_{N\rightarrow+\infty}\prod_{n=0}^{N-1}\frac{\mathscr{O}(X_{\tau_{n+1}})}{\mathscr{O}(X_{\tau_{n}})}$
$=\lim_{N\rightarrow+\infty}\prod_{n=0}^{N-1}\frac{\mathscr{O}(X_{\tau_{n}})+\mathscr{V}(\mathscr{O}){\rm d}\tau+o({\rm d}\tau^{2})}{\mathscr{O}(X_{\tau_{n}})}$
$=\exp\left[\int_{\tau_{I}}^{\tau_{F}}\mathscr{O}^{-1}\mathscr{V}(\mathscr{O}){\rm d}\tau\right]$
$=\exp\left[\int_{t_{I}}^{t_{F}}\gamma^{-1}\mathscr{O}^{-1}\mathscr{V}(\mathscr{O}){\rm d}t\right],$ (137)
where $\mathscr{V}$ is the tangent vector along $Y_{t}$, as presented by eq. (124) and/or eq. (132). It is worth noting that in this appendix, we will use $(y,k)$ to represent the coordinates of the tangent bundle, rather than $(x,p)$. This choice does not imply a coordinate transformation. It simply emphasizes that the discretized time corresponds to Alice’s proper time $t$.
Eq. (B) can be used to evaluate the continuum limit of certain ratios which are useful in proving the fluctuation theorem. For instance,
$\frac{k_{0}|_{Y_{N}}}{k_{0}|_{Y_{0}}}$ $=\exp\left[\int_{t_{I}}^{t_{F}}\frac{\gamma^{-1}}{k_{0}}\left\{\frac{k^{\mu}}{m}\frac{\partial}{\partial y^{\mu}}k_{0}+m\mathcal{A}^{\mu}\frac{\partial}{\partial k^{\mu}}k_{0}\right\}{\rm d}t\right]$
$=\exp\left[\int_{t_{I}}^{t_{F}}\frac{\gamma^{-1}}{k_{0}}\left\{\frac{1}{m}\partial_{\mu}g_{0\nu}k^{\mu}k^{\nu}+m\mathcal{A}_{0}\right\}{\rm d}t\right]$
$=\exp\left[\int_{t_{I}}^{t_{F}}\frac{\gamma^{-1}}{k_{0}}\left\{\frac{1}{2m}\partial_{0}g_{\mu\nu}k^{\mu}k^{\nu}+ma_{0}\right\}{\rm d}t\right],$ (138)
$\frac{g|_{Y_{0}}}{g|_{Y_{N}}}$ $=\exp\left[-\int_{t_{I}}^{t_{F}}\gamma^{-1}g^{-1}\frac{k^{\mu}}{m}\frac{\partial}{\partial y^{\mu}}g{\rm d}t\right],$ (139)
and
$\frac{\partial_{0}t|_{Y_{N}}}{\partial_{0}t|_{Y_{0}}}$ $=\exp\left[\int_{t_{I}}^{t_{F}}\frac{\gamma^{-1}}{\partial_{0}t}\frac{k^{\mu}}{m}\frac{\partial}{\partial\breve{y}^{\mu}}\frac{\partial}{\partial\breve{y}^{0}}t{\rm d}t\right]$
$=\exp\left[\int_{t_{I}}^{t_{F}}\frac{1}{\partial_{0}t}\left\{\frac{\partial}{\partial\breve{y}^{0}}\left(\gamma^{-1}\frac{k^{\mu}}{m}\partial_{\mu}t\right)-\partial_{\mu}t\frac{\partial}{\partial\breve{y}^{0}}\left(\gamma^{-1}\frac{k^{\mu}}{m}\right)\right\}{\rm d}t\right]$
$=\exp\left[\int_{t_{I}}^{t_{F}}\frac{1}{\partial_{0}t}\left\{\gamma^{-2}\frac{k^{\mu}}{m}\partial_{\mu}t\frac{\partial}{\partial\breve{y}^{0}}\gamma-\frac{\gamma^{-1}}{m}\partial_{0}t\frac{\partial}{\partial\breve{y}^{0}}k^{0}\right\}{\rm d}t\right]$
$=\exp\left[\int_{t_{I}}^{t_{F}}\gamma^{-1}\frac{\partial}{\partial t}\gamma{\rm d}t\right]\exp\left[\int_{t_{I}}^{t_{F}}\frac{\gamma^{-1}}{k_{0}}\frac{1}{2m}\partial_{0}g_{\mu\nu}k^{\mu}k^{\nu}{\rm d}t\right].$ (140)
Eq. (101) can be rewritten as
$\tilde{y}^{i}_{n+1}=h^{i}(\tilde{Y}_{n+1},\tilde{Y}_{n})=\frac{\tilde{k}^{i}_{\bar{n}}}{m}\gamma^{-1}|_{\tilde{Y}_{\bar{n}}}{\rm d}t+\tilde{y}^{i}_{n},$ (141)
which implies that the realization of $\tilde{y}^{i}_{n+1}$ is determined by the realization of $\tilde{Y}_{n}$ and $\tilde{k}^{i}_{n+1}$. The probability distribution of a random variable $\tilde{x}$ obeying the constraint $\tilde{x}=f(\tilde{x})$ is given by
$\Pr[\tilde{x}=x]=(1-f^{\prime}(x))\delta(x-f(x)).$ (142)
The generalization of this equation to general dimension is straightforward, telling that the probability distribution of $\tilde{y}^{i}_{n+1}$ under the conditions $\tilde{k}^{i}_{n+1}={k}^{i}_{n+1},\tilde{Y}_{n}=Y_{n}$ is given by
$\Pr[\tilde{y}_{n+1}=y_{n+1}|\tilde{k}_{n+1}={k}_{n+1},\tilde{Y}_{n}=Y_{n}]$
$\qquad=\det\left[\delta^{i}{}_{j}-\frac{\partial h^{i}}{\partial\hat{y}_{n+1}^{j}}\right]\left|\frac{\partial_{0}t}{\lambda\sqrt{g}}\right|_{y_{n+1}}\delta^{d}(y_{n+1}-h({Y_{n+1},Y_{n}}))$
$\qquad=\left[1+\gamma^{-2}\frac{k^{i}}{2m}\frac{\partial}{\partial\hat{{y}}^{i}}\gamma{\rm d}t+o({\rm d}t^{2})\right]_{Y_{\bar{n}}}\left|\frac{\partial_{0}t}{\lambda\sqrt{g}}\right|_{y_{n+1}}\Delta(Y_{n+1},Y_{n}),$ (143)
where $y_{n}$ denotes the sequence $(y_{n}^{1},\cdots,y_{n}^{d})$ and we will also use the notation $k_{n}=(k^{1}_{n},\cdots,k_{n}^{d})$. $\Delta(Y_{n+1},Y_{n})$ is defined as $\Delta(Y_{n+1},Y_{n}):=\delta^{d}(y_{n+1}-h(Y_{n+1},Y_{n}))$, which has the property
$\Delta(Y_{n+1},Y_{n})=\Delta(I(Y_{n}),I(Y_{n+1})).$ (144)
The appearance of $\partial/\partial\hat{y}^{i}$ in the second line of eq. (B) is due to the fact that $\tilde{y}^{i}_{n+1}$ only takes values in the configuration space, and its probability density is a scalar field on $\mathcal{S}_{n}$. The factor $|\partial_{0}t/(\lambda\sqrt{g})|$ arises from the change of coordinate volume element into invariant volume element, as indicated by eq. (2).
The single step transition probability can be rewritten as
$\Pr[\tilde{Y}_{n+1}|\tilde{Y}_{n}]=\Pr[\tilde{y}_{n+1}|\tilde{k}_{n+1},\tilde{Y}_{n}]\Pr[\tilde{k}_{n+1}|\tilde{Y}_{n}],$ (145)
where the second factor remains to be evaluated. Defining the function
${\rm d}W^{\mathfrak{a}}(Y_{n+1},Y_{n})$ :=(^$:=(\hat{\mathcal{R}}^{-1})^{\mathfrak{a}}{}_{i}|_{Y_{\bar{n}}}\left[k^{i}_{n+1}-k^{i}_{n}-F^{i}|_{Y_{\bar{n}}}{\rm d}t-\bar{F}^{i}|_{Y_{\bar{n}}}{\rm d}t\right]$
=(^$=(\hat{\mathcal{R}}^{-1})^{\mathfrak{a}}{}_{i}|_{Y_{\bar{n}}}\left[\gamma^{-1}m\mathcal{A}^{i}_{n}-F^{i}-\bar{F}^{i}\right]_{Y_{\bar{n}}}{\rm d}t,$ (146)
where $\mathcal{A}^{i}_{n}:=\gamma|_{Y_{\bar{n}}}(k^{i}_{n+1}-k^{i}_{n})/(m{\rm d}t)$ is the coordinate acceleration, then eq. (102) can be rewritten as
${\rm d}\tilde{W}^{\mathfrak{a}}_{n}={\rm d}W^{\mathfrak{a}}(\tilde{Y}_{n+1},\tilde{Y}_{n}).$ (147)
In the case that the realization of $\tilde{Y}_{n}$ is given, $\tilde{y}_{n+1}$ is determined by $\tilde{k}_{n+1}$. Therefore, eq. (147) gives the relation between ${\rm d}\tilde{W}^{\mathfrak{a}}_{n}$ and $\tilde{k}^{i}_{n+1}$, which can be used to calculate the conditional probability
$\Pr[\tilde{k}_{n+1}=k_{n+1}|\tilde{Y}_{n}=Y_{n}]=\left|\frac{k_{0}}{m\sqrt{g}}\right|_{Y_{n+1}}\det\left[T^{\mathfrak{a}}{}_{i}(Y_{n+1},Y_{n})\right]\Pr[{\rm d}\tilde{W}^{\mathfrak{a}}_{n}={\rm d}W^{\mathfrak{a}}(Y_{n+1},Y_{n})],$ (148)
where
$T^{\mathfrak{a}}{}_{i}(Y_{n+1},Y_{n}):=\frac{\partial}{\partial\breve{k}^{i}_{n+1}}{\rm d}W^{\mathfrak{a}}(Y_{n+1},Y_{n}),$ (149)
and $y_{n+1}$ appearing on the right hand side is regarded as a function $y_{n+1}(k_{n+1},Y_{n})$ determined by eq. (141). Using the implicit relationship (141) between $y_{n+1}$ and $k_{n+1}$, we can get the following relation by use of a differentiation with respect to $k_{n+1}^{i}$,
$\left[\delta^{i}{}_{\ell}-\frac{\partial\gamma^{-1}}{\partial\hat{{y}}_{n+1}^{\ell}}{\rm d}t\right]\frac{\partial y_{n+1}^{\ell}}{\partial\breve{k}_{n+1}^{j}}=\left[\frac{1}{2m}\delta^{i}{}_{j}\gamma^{-1}|_{Y_{\bar{n}}}+\frac{k^{i}_{\bar{n}}}{2m}\frac{\partial\gamma^{-1}}{\partial\breve{{k}}^{j}_{n+1}}\right]{\rm d}t,$ (150)
which indicates that $\partial y^{\ell}_{n+1}/\partial\breve{k}^{j}_{n+1}\sim o({\rm d}t)$. Therefore,
$T^{\mathfrak{a}}{}_{i}({Y_{n+1},Y_{n}})=\frac{\partial{\rm d}W^{\mathfrak{a}}}{\partial\breve{k}^{i}_{n+1}}+\frac{\partial y^{k}_{n+1}}{\partial\breve{k}^{i}_{n+1}}\frac{\partial{\rm d}W^{\mathfrak{a}}}{\partial\hat{y}^{k}_{n+1}}=\frac{\partial{\rm d}W^{\mathfrak{a}}}{\partial\breve{k}^{i}_{n+1}}+o({\rm d}t^{2})$
=(^$\qquad=(\hat{\mathcal{R}}^{-1})^{\mathfrak{a}}{}_{i}|_{Y_{\bar{n}}}+\frac{1}{2}\left\{m\gamma^{-1}\frac{\partial}{\partial\breve{k}^{i}}(\hat{\mathcal{R}}^{-1})^{\mathfrak{a}}{}_{j}\mathcal{A}^{j}_{n}-\frac{\partial}{\partial\breve{k}^{i}}[(\hat{\mathcal{R}}^{-1})^{\mathfrak{a}}{}_{j}F^{j}]-\frac{\partial}{\partial\breve{k}^{i}}[(\hat{\mathcal{R}}^{-1})^{\mathfrak{a}}{}_{j}\bar{F}^{j}]\right\}_{Y_{\bar{n}}}{\rm d}t,$ (151)
where terms of order $o({\rm d}t^{2})$ and higher have been omitted. Finally, the transition probability of the forward process can be expressed as
$\Pr[\tilde{Y}_{n+1}=Y_{n+1}|\tilde{Y}_{n}=Y_{n}]$
$\qquad=\left[1+\frac{k^{i}}{2m}\gamma^{-2}\frac{\partial}{\partial\hat{y}^{i}}\gamma{\rm d}t+o({\rm d}t^{2})\right]_{Y_{\bar{n}}}\left|\frac{k_{0}\partial_{0}t}{m\lambda g}\right|_{Y_{n+1}}\Delta(Y_{n+1},Y_{n})\det\left[T^{a}{}_{i}(Y_{n+1},Y_{n})\right]$
×$\qquad\quad\times\Pr[{\rm d}\tilde{W}^{\mathfrak{a}}_{n}={\rm d}W^{\mathfrak{a}}{(Y_{n+1},Y_{n})}].$ (152)
The single step transition probability in the reversed process can be evaluated following a similar procedure, yielding
$\Pr[\tilde{Y}^{-}_{N-n}=I(Y_{n})|\tilde{Y}^{-}_{N-n-1}=I(Y_{n+1})]$
$\qquad=\left[1+\frac{k^{i}}{2m}\gamma^{-2}\frac{\partial}{\partial\hat{{y}}^{i}}\gamma{\rm d}t+o({\rm d}t^{2})\right]_{I(Y_{\bar{n}})}\left|\frac{k_{0}\partial_{0}t}{m\lambda g}\right|_{I(Y_{n})}$
×$\qquad\quad\times\Delta(I(Y_{n}),I(Y_{n+1}))\det\left[T^{a}{}_{i}(I(Y_{n}),I(Y_{n+1}))\right]\Pr[{\rm d}\tilde{W}^{\mathfrak{a}}_{n}={\rm d}W^{\mathfrak{a}}(I(Y_{n}),I(Y_{n+1}))]$
$\qquad=\left[1-\frac{k^{i}}{2m}\gamma^{-2}\frac{\partial}{\partial\hat{y}^{i}}\gamma{\rm d}t+o({\rm d}t^{2})\right]_{Y_{\bar{n}}}\left|\frac{k_{0}\partial_{0}t}{m\lambda g}\right|_{Y_{n}}\Delta(Y_{n+1},Y_{n})\det\left[T^{a}{}_{i}(I(Y_{n}),I(Y_{n+1}))\right]$
×$\qquad\quad\times\Pr[{\rm d}\tilde{W}^{\mathfrak{a}}_{n}={\rm d}W^{\mathfrak{a}}(I(Y_{n}),I(Y_{n+1}))].$ (153)
The Jacobian matrix $T^{a}{}_{i}(I(Y_{n}),I(Y_{n+1}))$ in the reversed process reads
$T^{\mathfrak{a}}{}_{i}(I(Y_{n}),I(Y_{n+1}))$
=(^$\qquad=(\hat{\mathcal{R}}^{-1})^{\mathfrak{a}}{}_{i}|_{I(Y_{\bar{n}})}+\frac{1}{2}\left\{m\gamma^{-1}\frac{\partial}{\partial\breve{k}^{i}}(\hat{\mathcal{R}}^{-1})^{\mathfrak{a}}{}_{j}\mathcal{A}^{j}_{n}-\frac{\partial}{\partial\breve{k}^{i}}[(\hat{\mathcal{R}}^{-1})^{\mathfrak{a}}{}_{j}F^{j}]-\frac{\partial}{\partial\breve{k}^{i}}[(\hat{\mathcal{R}}^{-1})^{\mathfrak{a}}{}_{j}\bar{F}^{j}]\right\}_{I(Y_{\bar{n}})}{\rm d}t$
=(^$\qquad=(\hat{\mathcal{R}}^{-1})^{\mathfrak{a}}{}_{i}|_{Y_{\bar{n}}}+\frac{1}{2}\left\{-m\gamma^{-1}\frac{\partial}{\partial\breve{k}^{i}}(\hat{\mathcal{R}}^{-1})^{\mathfrak{a}}{}_{j}\mathcal{A}^{j}_{n}+\frac{\partial}{\partial\breve{k}^{i}}[(\hat{\mathcal{R}}^{-1})^{\mathfrak{a}}{}_{j}F^{j}]-\frac{\partial}{\partial\breve{k}^{i}}[(\hat{\mathcal{R}}^{-1})^{\mathfrak{a}}{}_{j}\bar{F}^{j}]\right\}_{Y_{\bar{n}}}{\rm d}t.$ (154)
In the continuum limit, we have
$\lim_{N\rightarrow+\infty}\prod_{n=0}^{N-1}\frac{\det[T^{a}{}_{i}(Y_{n+1},Y_{n+1})]}{\det[T^{a}{}_{i}(I(Y_{n}),I(Y_{n+1}))]}$
$\qquad=\exp\left[\int_{t_{I}}^{t_{F}}\hat{\mathcal{R}}^{i}{}_{\mathfrak{a}}\left\{m\gamma^{-1}\frac{\partial}{\partial\breve{k}^{i}}(\hat{\mathcal{R}}^{-1})^{\mathfrak{a}}{}{j}\mathcal{A}^{j}-\frac{\partial}{\partial\breve{k}^{i}}[(\hat{\mathcal{R}}^{-1})^{\mathfrak{a}}{}_{j}\bar{F}^{j}]\right\}{\rm d}t\right]$
$\qquad=\exp\left[\int_{t_{I}}^{t_{F}}\gamma^{-1}\hat{\mathcal{R}}^{i}{}_{\mathfrak{a}}\frac{\partial}{\partial\breve{k}^{i}}(\hat{\mathcal{R}}^{-1})^{\mathfrak{a}}{}_{j}[ma^{j}-\mathcal{F}^{j}_{\text{em}}]{\rm d}t\right]\exp\left[-\int_{t_{I}}^{t_{F}}\frac{\partial}{\partial\breve{k}^{i}}F^{i}{\rm d}t\right],$ (155)
where the exponent in the last term can be expanded as
$-\frac{\partial}{\partial\breve{k}^{i}}F^{i}=$ $-\left(\mathcal{F}^{i}_{\text{em}}-\frac{1}{m}\varGamma^{i}{}_{\alpha\beta}k^{\alpha}k^{\beta}\right)\frac{\partial}{\partial\breve{k}^{i}}\gamma^{-1}+\gamma^{-1}\frac{k^{\mu}}{m}g^{-1}\frac{\partial}{\partial y^{\mu}}g$
$-\frac{\gamma^{-1}}{k_{0}}\left[\frac{1}{m}\partial_{0}g_{\alpha\beta}k^{\alpha}k^{\beta}+(\mathcal{F}_{\text{em}})_{0}\right].$ (156)
Meanwhile, we also have
$\lim_{N\rightarrow+\infty}\prod_{n=0}^{N-1}\frac{\left[1+\frac{k^{i}}{2m}\gamma^{-2}\frac{\partial}{\partial\hat{y}^{i}}\gamma{\rm d}t+o({\rm d}t^{2})\right]_{Y_{\bar{n}}}}{\left[1-\frac{k^{i}}{2m}\gamma^{-2}\frac{\partial}{\partial\hat{y}^{i}}\gamma{\rm d}t+o({\rm d}t^{2})\right]_{Y_{\bar{n}}}}=\exp\left[\int_{t_{I}}^{t_{F}}\gamma^{-2}\frac{k^{i}}{m}\frac{\partial}{\partial\hat{y}^{i}}\gamma{\rm d}t\right].$ (157)
Moreover, the continuum limit of the ratio of the probabilities of the Gaussian increments reads
$\lim_{N\rightarrow+\infty}\prod_{n=0}^{N-1}\frac{\Pr[{\rm d}\tilde{W}^{\mathfrak{a}}_{n}={\rm d}W^{\mathfrak{a}}(Y_{n+1},Y_{n})]}{\Pr[{\rm d}\tilde{W}^{\mathfrak{a}}_{n}={\rm d}W^{\mathfrak{a}}(I(Y_{n}),I(Y_{n+1}))]}=\exp\left[\int_{t_{I}}^{t_{F}}2\gamma^{-2}(\hat{\mathcal{D}}^{-1})_{ij}(m\mathcal{A}^{i}-\gamma F^{i})\gamma\bar{F}^{j}{\rm d}t\right],$ (158)
where the integrads can also be expanded as
$2\gamma^{-2}(\hat{\mathcal{D}}^{-1})_{ij}(m\mathcal{A}^{i}-\gamma F^{i})\gamma\bar{F}^{j}$
$\qquad=2\gamma^{-1}(\mathcal{D}^{-1})_{ij}(ma^{i}-\mathcal{F}^{i}_{\text{em}})\left(\mathcal{F}^{j}_{\text{dp}}+\mathcal{F}^{j}_{\text{add}}-\frac{1}{2}\mathcal{D}^{jk}\gamma^{1/2}\nabla_{k}^{(h)}\gamma^{-1/2}\right)$
$\qquad=\gamma^{-1}\frac{1}{T_{\rm B}}(ma^{\mu}-\mathcal{F}^{\mu}_{\text{em}})U_{\mu}+\gamma^{-1}(\mathcal{R}^{-1})^{a}_{j}\nabla_{i}^{(h)}\mathcal{R}^{i}{}_{a}(ma^{j}-\mathcal{F}^{j}_{\text{em}})$
$\qquad\quad+\frac{1}{2}\gamma^{-2}(ma^{i}-\mathcal{F}^{i}_{\text{em}})\nabla_{i}^{(h)}\gamma$
$\qquad=\gamma^{-1}\frac{1}{T_{\rm B}}(ma^{\mu}-\mathcal{F}^{\mu}_{\text{em}})U_{\mu}+\gamma^{-1}(\hat{\mathcal{R}}^{-1})^{a}{}_{j}\frac{\partial}{\partial\breve{k}^{i}}\hat{\mathcal{R}}^{i}{}_{a}(ma^{j}-\mathcal{F}^{j}_{\text{em}})$
$\qquad\quad+\gamma^{-2}(ma^{i}-\mathcal{F}^{i}_{\text{em}})\frac{\partial}{\partial\breve{k}^{i}}\gamma-\frac{\gamma^{-1}}{k_{0}}(ma_{0}-(\mathcal{F}_{\text{em}})_{0}).$ (159)
In the end, by combining all the continuum limit together, we obtain the ratio of the conditional probabilities for the forward and reversed processes,
$\lim_{N\rightarrow+\infty}\prod_{n=0}^{N-1}\frac{\Pr[\tilde{Y}_{n+1}=Y_{n+1}|\tilde{Y}_{n}=Y_{n}]}{\Pr[\tilde{Y}^{-}_{N-n}=I(Y_{n})|\tilde{Y}^{-}_{N-n-1}=I(Y_{n+1})]}$
[…]
$\qquad=\frac{\lambda|_{Y_{0}}}{\lambda|_{Y_{N}}}\exp\left[\int_{t_{I}}^{t_{F}}\gamma^{-1}\frac{1}{T_{\rm B}}(ma^{\mu}-\mathcal{F}^{\mu}_{\text{em}})U_{\mu}{\rm d}t\right]$
×$\qquad\quad\times\exp\left[\int_{t_{I}}^{t_{F}}\gamma^{-2}\left(\frac{k^{i}}{m}\frac{\partial}{\partial\hat{y}^{i}}\gamma+\gamma\frac{\partial}{\partial t}\gamma+m\mathcal{A}^{i}\frac{\partial}{\partial\breve{k}^{i}}\gamma\right){\rm d}t\right]$
$\qquad=\frac{\lambda|_{Y_{0}}}{\lambda|_{Y_{N}}}\exp\left[\int_{t_{I}}^{t_{F}}\gamma^{-1}\frac{1}{T_{\rm B}}(ma^{\mu}-\mathcal{F}^{\mu}_{\text{em}})U_{\mu}{\rm d}t\right]\exp\left[\int_{t_{I}}^{t_{F}}\gamma^{-2}\mathscr{V}(\gamma){\rm d}t\right]$
$\qquad=\frac{(\lambda\gamma^{-1})|_{Y_{0}}}{(\lambda\gamma^{-1})|_{Y_{N}}}\exp\left[\int_{t_{I}}^{t_{F}}\gamma^{-1}\frac{1}{T_{\rm B}}(ma^{\mu}-\mathcal{F}^{\mu}_{\text{em}})U_{\mu}{\rm d}t\right],$ (160)
which is exactly eq. (5) appeared in the main text.